### v2

# Inheritune: Training Smaller Yet More Attentive Language Models

Abstract
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, primarily due to the transformer architecture and its self-attention mechanism. However, we observe that in standard decoder-style LLMs, attention matrices degenerate to single-column for deeper layers. Layers in this state are unable to learn anything meaningful and mostly redundant; we refer to these as lazy layers. The goal of this paper is to train smaller models by eliminating this structural inefficiency without compromising performance.
Motivated by this observation, we propose Inheritune, a simple yet effective training recipe for developing smaller, high-performing language models. Smaller models trained with $\mathsf{Inheritune}$ inherit early transformer layers from a larger pre-trained model, then retrain and progressively expand until they match or exceed the performance of the larger model. We demonstrate that $\mathsf{Inheritune}$ enables the training of various sizes of GPT-2 models on datasets like OpenWebText-9B and FineWeb_Edu. Models trained with $\mathsf{Inheritune}$, despite having significantly fewer layers, match or even surpass the performance of their larger counterparts. For instance, our 16-layer GPT-2 medium variant achieves comparable performance to the standard 24-layer GPT-2 medium model. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked together, building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration, caused by a collapse in the attention rank Noci et al. (2022); Dong et al. (2021); He et al. (2023). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e., their mass is concentrated in a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerate, as lazy layers.
Motivated by this new finding, we aim to develop performant small base language models (LMs) that utilize weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple, yet effective, training approach as $\mathsf{Inheritune}$.
(a) Rank analysis of vanilla GPT-2 Medium
(b) Matrix mass analysis of vanilla GPT-2 Medium
(c) Model initialization with later layers
(d) Rank analysis of vanilla GPT-2 Large
(e) Matrix mass analysis of vanilla GPT-2 Large
(f) Model initialization with later layers
Figure 1: Attention matrices of many deeper layers often degenerate to single-column matrices in regular decoder style LLMs, layers with fully degenerate attention fail to learn meaningful representations. We computed a single attention matrix with 100 tokens from the OpenWebText validation set with 4M tokens. Next we performed 100 runs and plotted the mean and std of the max rank and mass as a function of layers for our rank and mass analysis respectively. Figure (a and d): An analysis of a 24-layer GPT-2 medium and a 36-layer GPT-2 large shows the max rank of the attention matrices across all layers. Figure (b and e): A closer look at the same GPT-2 models also reveals that the dominant mass proportion of several attention matrices is concentrated in a single-column, particularly in deeper layers. Figure (c and f): When initializing 12-layer and 18-layer variants of the vanilla GPT-2 medium and GPT-2 large models with deeper layers (Lazy layers) exhibiting degenerated attention–their performance is comparable to models with random initialization. However, initializing models with early layers leads to significantly better generalization and convergence.
In summary, our key contributions are as follows:
  1. 1.
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (see Figure 1 and Figure 2).
  2. 2.
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText dataset with 9B tokens (with data repetition) and the FineWeb_edu dataset with 100B tokens (without data repetition).
  3. 3.
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Table 1), model initialization and efficient training baselines (refer Table 2), and models trained using two knowledge distillation techniques (refer Figure 3). In settings where training tokens are not repeated, we observe similar trends (refer Figure 4).
L4 H2
L7 H3
L20 H2
L22 H4
24 layer GPT2 medium model trained from scratch.
L4 H2
L7 H3
L11 H2
L15 H1
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
Figure 2: Inheritune preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected strings, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s later layers.
2 Attention Degeneration in Standard Decoder-Style LLMs
Preliminaries:
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times d}$, where $T$ denotes the sequence length (number of tokens), and $d$ represents the embedding dimension or hidden size of the model. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times d}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weigh the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as
$\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$ (1)
where the queries $Q=XW_{Q}$,  keys $K=XW_{K}$,  and values $V=XW_{V}$ are linear transformations of the input $X$. Here, $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ are the weight matrices for the queries, keys, and values, respectively. Typically, $d_{k}=d_{v}=\frac{d}{h}$, where $h$ is the number of attention heads. In this single-head scenario, we set $d_{k}=d_{v}=d$.
The attention matrix $A(X)\in\mathbb{R}^{T\times T}$ captures the pairwise attention scores between all token positions in the sequence. The softmax is applied row-wise. The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors.
Previous research by Dong et al. (2021) and He et al. (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model. This phenomenon, known as the rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly. Noci et al. (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$. Interestingly, they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers, which affects the overall trainability of the transformer-based models. However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layer norms, and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
Approximate Rank Computation of Attention Matrices
In this paper, we deeply analyzed the structure of attention matrices to diagnose the presence of rank collapse or similar phenomena in standard transformer-based LLMs using GPT-2 models. For our first analysis, we compute the approximate rank (referred to as rank hereafter) of $A(X)$ for all attention heads within each layer. Formally, we began by computing the Singular Value Decomposition (SVD) of $A(X)=U\Sigma V^{\top}$, where the diagonal entries $\{\sigma_{i}\}_{i=1}^{T}$ of $\Sigma$ represent the singular values, quantifying the variance captured by each corresponding singular vector of $A(X)$. To determine the minimal number of singular values required to capture 90% of the total variance, we solved: $k^{*}=\min\left\{k\in\{1,2,\ldots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq 0.90\right\}$. Here, $k^{*}$ is the approximate rank of $A(X)$ computed using the explained variance method.
Dominant Single-Column Structure in Attention Matrices
We further investigated the dominant structure of these rank-1 attention matrices and observed that, on average, many of these matrices have their mass concentrated in a single column. This intrinsic structure can be viewed as a special case of rank-1 attention matrices. To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$. Next, to determine the minimal number of columns required to capture 90% of the total mass, we solved; $m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq 0.90\right\}$. Here, $m^{*}$ represents the minimal number of columns needed for the cumulative column mass ratios to reach or exceed 90%.
Degeneration of Attention Matrices in GPT-2 Models
In Figure 1, we present the layer-wise analysis of rank and mass. For this analysis, we computed $A(X)$ using 100 randomly selected samples from the validation set of OpenWebText valset with 4M tokens, each with a sequence length of $T=100$ tokens, across all attention heads within each layer. Next, we computed the rank with a 90% variance threshold, and for every layer, we chose the maximum rank across all the heads. In Figure 1(a) and 1(d) we plotted maximum rank as a function of layers for 100 runs with mean and standard deviation (std); it is quite evident that many deeper layers exhibit all rank-1 attention matrices. We highlight that this rank collapse occurs for both the GPT-2 medium and GPT-2 Large models with skip connections and FFNs, extending the findings of Dong et al. (2021) and Noci et al. (2022) to standard LLM architectures. Next, based on our mass analysis, we demonstrate that most of the rank collapsed attention matrices are also single-column matrices as depicted in Figure 1(b) and Figure 1(e). We follow the analysis protocol as discussed in the rank analysis. This degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit complete degeneration of all attention matrices across all attention heads, indicating reduced performance and less effective token mixing.
How much transferable knowledge do lazy layers hold compared to their early counterparts? To answer this question, we initialized a 12-layer GPT-2 medium variant [^1] A variant shares the same configurations as the parent model but has fewer layers. and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. These pre-trained models are trained on the OpenWebText-9B dataset for 100K steps. We then fine-tuned these GPT-2 variants on the same dataset for an additional 10K steps. For comparison, we conducted two baseline experiments in which the GPT-2 variants were initialized either with the first few transformer blocks or with random initialization. As shown in Figures 1(c) and 1(f), models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provides additional evidence that lazy layers with fully degenerate attention fail to learn meaningful representations.
2.1 Attention Pattern Visualization
To provide further evidence of lazy layers and provide a preview of our solution, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 2 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. We visualized full attention and not causal attention for the sake of better visualization. In practice, GPT-2 models compute causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over three runs.
Algorithm 1 Inheritune: Training Recipe for Small Language Models
1:Reference model $\mathcal{M}_{\text{ref}}$ with $k$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
2:Initialize $\mathcal{M}_{\text{tgt}}$ with first $n=k/2$ layers from $\mathcal{M}_{\text{ref}}$
3:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
4:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
5:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
6:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
7:end while
8:return Optimized model $\mathcal{M}_{\text{tgt}}$
In the 24-layer model trained from scratch (top row of Fig. 2), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes the model more attentive and addresses attention degeneration, potentially leading to more efficient models in a compact size (also refer to Figure 9). We will discuss these results in more detail after introducing our method, but this preview underscores the promise of our approach.
3 Inheritune: Our Proposed Training Recipe
This section offers a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
Recall that we have established the issue of attention degeneration with two motivating examples, highlighting specific inefficiencies in pre-trained LLMs. In this paper, we turn this challenge into an opportunity to create smaller base LMs that are equally performant, achieving similar or lower validation loss compared to their larger, less efficient counterparts, which we refer to as reference models. Our proposed solution builds on two key insights: (1) the early layers of deep LLMs provide effective model initialization, and (2) multiple lazy layers can be collapsed into fewer layers and re-trained to improve the model capacity.
Setup:
We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$. Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps. We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
We now present $\mathsf{Inheritune}$, our proposed training recipe for efficiently developing small base language models (LMs). $\mathsf{Inheritune}$ operates on the principle of zero-shot initialization and progressive growth. The $\mathsf{Inheritune}$ process consists of three main steps, which we present below and formalize in Algorithm 1:
  1. 1.
    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $n=k/2$ layers of $\mathcal{M}_{\text{ref}}$, including weights, prediction head, and token embeddings.
  2. 2.
    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
  3. 3.
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size and repeat steps 1-2 until the desired performance is achieved.
With our method now formally described, we turn to empirical validation. In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes. In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing information on the adaptability of our approach.
4 Experiments
We evaluated $\mathsf{Inheritune}$ through a series of comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M) and GPT-2 medium (355M) models, Radford et al. (2019) pre-trained on the 9B tokens OpenWebText dataset (Gokaslan & Cohen, 2019). These models are trained with data repetition, meaning data is randomly sampled with replacement during batch creation. This experimental setup is adapted from Liu et al. (2023); Sanyal et al. (2024). For evaluation, we compare model(s) trained with $\mathsf{Inheritune}$ with baselines from three key settings: a) baseline models trained from scratch with random initialization, b) baseline models trained using various zero-shot initialization techniques, and c) baseline models trained with knowledge distillation. Table 10 provides detailed specifications for all models used in our experiments. Finally, we conduct a thorough ablation study of our initialization strategy, focusing on the 16-layer GPT-2 medium variants.
We provide experimental details of our $\mathsf{Inheritune}$ training recipe using a GPT-2 large model as an example; a similar procedure was applied to train other models. Our methodology for applying $\mathsf{Inheritune}$involves the following steps:
  1. 1.
    Reference Model: We train the full 36-layer GPT-2 Large model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
  2. 2.
    Model initialization We initialize an 18-layer model ($n=k/2$) using the trained 36-layer model as a reference.
  3. 3.
    Training and Evaluation: We train the 18-layer model on $\mathcal{D}_{\text{train}}$ for $T$ steps and evaluate its validation loss.
  4. 4.
    Iterative Refinement: If the smaller model’s performance is inferior, we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
Models Layers Initialization Steps Pre-train Downstream (0-shot)
Val loss ($\downarrow$) Wikitext ($\downarrow$) Lambada
GPT-2 Medium 24 rand init 100K 2.81 31.93 36.54
16 rand init 100K 2.86 33.67 34.60
16 rand init 200K 2.83
12 Ours 100K 2.87
14 Ours 100K 2.84
Final Model $\longrightarrow$ 16 Ours 100K 2.81 32.04 35.96
GPT-2 Large 36 rand init 100K 2.85 34.84 34.14
18 rand init 100K 2.97 37.63 30.97
18 rand init 200K 2.84
18 Ours 100K 2.80 35.38 34.64
GPT-2 xLarge 48 rand init 100K 2.65 25.45 39.90
24 rand init 100K 2.69 28.32 38.46
24 rand init 200K 2.62
24 Ours 100K 2.64 25.52 43.30
Table 1: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines. Metrics include pre-training validation loss ($\downarrow$), zero-shot Wikitext ($\downarrow$) and Lambada performance. Note: GPT-2 xLarge and Large took one round of training; GPT-2 Medium took three rounds.
Baselines trained from scratch (rand init.) :
We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
Baselines trained with various model initialization and efficient training techniques.
Here, we compare our model derived using $\mathsf{Inheritune}$ to models of similar size trained with various zero-shot model initialization and efficient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process to other models.
  1. 1.
    Stacking Gong et al. (2019); J. Reddi et al. (2023) is a model initialization and efficient (stage-wise) training recipe. We trained a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17. Finally, we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
  2. 2.
    Hybrid stacking: Hybrid stacking is stacking that utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights. We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to an 18-layer model by copying the weights to layers 0-17. We then trained this new 18-layer GPT-2 variant for 100K steps.
  3. 3.
    Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth. We copied the weights of the first half of the attention heads (0-9) and MLPs of the GPT-2 large reference model into the baseline GPT-2 variant with half the width but all layers.
Baselines trained with Knowledge Distillation
As a baseline, we first apply logit-based knowledge distillation (KD) Hinton et al. (2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly. For the second baseline, we use a DistillBERT-style approach Sanh et al. (2019), where the 0-11 layers of the student model are initialized with each alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher. Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model). Additional results with 12-layer GPT-2 medium variant(s) trained from scratch, using KD baseline and our recipe can be found in Figure 8.
Models Layers Recipe Steps Pre-train Val loss ($\downarrow$)
GPT-2 Medium 24 half-width 100K 3.04
16 stacking 100K 2.84
16 hybrid-stacking 100K 2.83
16 Ours 100K 2.81
GPT-2 Large 36 half-width 100K 3.06
18 stacking 100K 2.87
18 hybrid-stacking 100K 2.89
18 Ours 100K 2.80
GPT-2 xLarge 48 half-width 100K 2.77
24 stacking 100K 2.65
24 hybrid-stacking 100K 2.64
24 Ours 100K 2.64
Table 2: Inheritune outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
4.1 Results and Discussions
Models trained with Inheritune outperforms much larger models trained from scratch.
We present our main results in Table 1. Our 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium achieve comparable or lower validation losses than their full-sized counterparts when trained for the same number of steps (100K steps). Our GPT-2 xlarge and GPT-2 large variants undergo one round of $\mathsf{Inheritune}$ training, while for GPT-2 medium, we perform three rounds of training with 12, 14, and 16 layers. We also evaluated all models on two next-word prediction downstream tasks in a zero-shot setting using the Wikitext Merity et al. (2016) and Lambada Paperno et al. (2016) datasets. The downstream performance of our GPT-2 models derived using the $\mathsf{Inheritune}$ recipe matches their much larger reference models. From the convergence perspective, some prior works have made connections between over-parameterization and faster convergence Bengio et al. (2005); Vaswani et al. (2018). In supplementary Figure 5, we show that the small LMs derived with $\mathsf{Inheritune}$, although smaller compared to their reference models, still converge as fast as their large-size reference models.
Models trained with Inheritune outperform same-sized models trained from scratch.
Table 1 demonstrates that GPT-2 variants trained with $\mathsf{Inheritune}$ outperform their same-sized counterparts trained from scratch, both when trained for the same number of steps and even when trained for double the steps (200K). This result underscores the efficiency of our approach. The only exception is the 24-layer GPT-2 xlarge variant, which surpasses both our model and the full-size model when trained for 200K steps.
Models trained with Inheritune outperform all zero-shot model initialization and efficient training baselines.
In Table 2, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. Although stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. In all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 6.
Figure 3: A 16-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized model trained with Logit-based distillation baselines. We conducted vanilla KD Hinton et al. (2015) and DistillBERT-style KD Sanh et al. (2019) with teacher initialization, using a 24-layer GPT-2 medium as the teacher for both KD baselines.
Distillation vs Inheritune.
In Figure 3, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K steps. Additional distillation experiments can be found in the supplementary material.
4.2 Ablations
Layers Initialization Steps Val loss ($\downarrow$)
16 attn (w/ layernorm) 100K 2.84
16 mlp (w/ layernorm) 100K 2.85
16 attn+mlp (w/o layernorm) 100K 2.80
16 Ours (attn+mlp w/ layernorm) 100K 2.81
Table 3: Impact of initializing various sub-modules within a transformer block. We compare the validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText-9B dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 7.
We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence. For these ablations, we fixed the model to a 16-layer GPT-2 medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 medium reference model. We initialize the transformer blocks with 1) attention ((key, query, value, and projection) and the layernorm [^2] In GPT-2 models layernorm blocks are parameterized. weights (attn w/ layernorm), 2) attention and mlp weights without the layer-norm (attn+mlp w/o layernorm), and 3) mlp weights with the layer norm (mlp w/ layernorm). We emphasize that $\mathsf{Inheritune}$ performs initialization by inheriting attention and mlp weights with the layer norm (attn+mlp w/ layernorm).
As shown in Table 3, models trained with attention and mlp weights demonstrated the best performance, regardless of the layer norm initialization. A detailed validation loss vs training steps plot is presented in supplementary Figure 7. We conclude that initializing both attention and MLP weights provides a clear advantage. Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
5 Training without Data Repetition
(a) GPT2-Large
(b) GPT2-Medium
Figure 4: Models derived using Inheritune without data repetition converge faster and match the final validation loss of the full-sized model despite using fewer layers. Additionally, the model trained using $\mathsf{Inheritune}$ demonstrates data efficiency, achieving a lower validation loss in fewer steps compared to its full-sized and half-sized counterparts until 80% of the training.
Are the gains we observe due to the Inheritune recipe merely a consequence of over-fitting due to data repetition?
To investigate this, we conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023a); Biderman et al. (2023). Moreover, we utilized a high-quality pre-training dataset, Fineweb_edu Penedo et al. (2024), which contains 100B tokens and has been deduplicated and filtered to ensure high data quality.
We train a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following Algorithm 1. Finally, we trained the 16-layer baseline variants from scratch for comparison. All these models are trained for 100K steps. The model configurations and training hyper-parameters can be found in supplementary Table 10.
As shown in Figure 4, our GPT-2 variants trained using $\mathsf{Inheritune}$ consistently perform on par with their full-size counterparts and outperform their same-size counterparts in terms of training loss. In LLM pre-training literature, where data is not repeated, training loss has been shown to be a reliable metric Touvron et al. (2023a, b). Furthermore, we performed zero-shot downstream evaluations using lm-evaluation-harness Gao et al. (2024) on a variety of tasks, including ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), Hellaswag Zellers et al. (2019), and PIQA Bisk et al. (2020). As shown in Table 4, on average, the models demonstrate superior performance than the baseline models trained from scratch.
Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
(acc) (acc) (acc) (acc norm) (acc)
GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
Ours 16 51.26 61.81 73.8 30.55 23 48.08
GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
rand init 16 50.34 63.11 75 30.86 21.56 48.17
Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
Table 4: Models trained with Inheritune outperform both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
6 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied as a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation, i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggests that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to previous work, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023), which employs a stage-wise training strategy that uses weights from initial layers to initialize later layers, has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers-though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating a smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
7 Conclusion
In this study, we identified a structural flaw in the attention mechanism of deep decoder-style transformers, which are currently the de-facto architectural choice for LLMs. Specifically, we discovered that the attention matrices in many deeper layers tend to lose rank and converge into single-column matrices, resulting in what we term lazy layers. This phenomenon causes the model to lose attention and leads to structural inefficiencies in LLMs.
To avoid this issue, we propose a simple training recipe $\mathsf{Inheritune}$ where we inherit a few early blocks from a larger reference model and continue training and expanding the model with more blocks until it matches the performance of the reference model. We evaluated our method using 48-layer, 36-layer, and 24-layer GPT-2 xlarge, GPT-2 large, and GPT-2 medium reference models. Through this process, we successfully developed smaller models with 24, 18, and 16 layers without losing any performance compared to their larger counterparts while utilizing the OpenWebText-9B dataset. We observed thematically similar results with GPT-2 large and GPT-2 medium models trained on FineWeb_edu without data repetition. These findings offer a novel perspective on developing efficient small base language models from existing large pre-trained models, potentially democratizing LLM pre-training for a wider community.
Acknowledgments
We thank Nived Rajaraman, for his helpful feedback and discussion during this work.
Supplementary Materials
Contents
  • A: Supplementary Experiments
  • B: Additional Experiments on Low Data Regime
  • C: Implementation Details
Appendix A Supplementary Experiments
We provide additional training plots for our main results discussed in Section 4.1 as shown in Figure 5 and Figure 6. In Figure 5 (also refer Table 1 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 6 (also refer Table 2) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
In Figure 7, we present the training curves of models trained during ablation as discussed in Section 4.2.
Knowledge Distillation
Recall we have already discussed distillation as a baseline in Section 4.1 and associated Figure 3. We perform an additional experiment in the same setting i.e. knowledge distillation as a baseline. Here we trained GPT-2 medium variants with 12 layers (half the number of a vanilla GPT-2 medium). We trained three models. First we distilled a 24-layer GPT-2 medium (teacher) to a 12-layer GPT-2 medium variant (student) and this student is initialized with all the alternate layers of the teacher. This setting is exactly same as discussed in DistillBERT Sanh et al. (2019). Next we trained two GPT-2 medium variants one from scratch (vanilla training) and the other with $\mathsf{Inheritune}$ recipe. Model trained with our recipe beats model trained with distillation. We defer a through investigation of distillation compared to $\mathsf{Inheritune}$ to future work.
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 2. Following up on our previous discussions in Figure 9 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers (refer Figure 9) compared to it’s larger counterpart trained form scratch. We performed rank analysis utilizing 24-layer GPT-2 medium and our 16-layer GPT-2 model trained using $\mathsf{Inheritune}$.
Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al. (2022). The vanishing gradients is caused when the norm of the gradients Bengio et al. (1994) are so small that it fails to generate meaningful back-propagation signal. Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model trained from scratch, despite being smaller. Training GPT-2 xlarge, GPT-2 large and GPT-2 medium vanilla models from scratch and our variants with OpenWebText-9B for 100K steps.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 6: Models derived using Inheritune outperform three zero-shot initialization and efficient training baselines in terms of final validation loss. Our models demonstrate better convergence and generalization compared to all baselines. We trained GPT-2 xlarge, GPT-2 large and GPT-2 medium variants on OpenWebText-9B for 100K steps using baseline model initialization and efficient training techniques and our $\mathsf{Inheritune}$ training recipe.
Figure 7: Full training curves of 16-layer GPT-2 variants trained during ablations. We analyze $\mathsf{Inheritune}$ approach while initializing some specific sub-modules in transformer blocks. Here, we initialize each transformer block of a 16-layer GPT-2 medium variant with three different configurations. First, we separately initialize attention and MLPs (FFNs) submodules; second, we initialize the attention and MLP weights while randomly initializing the layer norms. Finally, we perform $\mathsf{Inheritune}$-initialize only the attention and MLP weights with all the respective layer norms.
Figure 8: A 12-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized models trained from scratch and with Logit-based distillation with teacher initialization baseline. Three 12-layer GPT-2 medium variants were trained: (1) a distilled model initialized with alternate layers from a 24-layer GPT-2 medium teacher, following the DistillBERT setup Sanh et al. (2019); (2) a model trained from scratch (vanilla training); and (3) a model trained using the $\mathsf{Inheritune}$ recipe. The model trained with $\mathsf{Inheritune}$ outperforms both the distillation-based model and the one trained from scratch, demonstrating the effectiveness of our approach.
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 9: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers, indicating rank collapse. (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach in preserving expressive power.
Appendix B Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
Figure 10: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding. We compare our model’s performance with reference model-OpenLLamA-3B (2x size), other small base LMs of size 1B-2B parameters such as MPT-1.3B, OPT-1.3B, Pythia-1.4B (pre-trained from scratch) and ShearLLaMA-1.5B (pruned and continually trained using existing large base LM).
In this section, we aim to investigate the efficacy of $\mathsf{Inheritune}$ in a data and compute-constrained setting. We train a 1.5B parameter small base LM with only 1B tokens using a 3B parameter base LM on a single GPU (A6000) for less than half a day.
We assume the existence of a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$. However, this full training data is unavailable, and we only have a random tiny subset $\hat{\mathcal{D}}_{\text{train}}\sim\mathcal{D}_{\text{train}}$. We use OpenLLaMA-3B version 1 as the reference model pre-trained with 1T tokens from the RedPajama V1 dataset, which contains data from various domains such as common crawl, C4, Wikipedia, books, arXiv papers, GitHub, and Stack Exchange. We take 1B randomly sampled tokens [^3] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample from the RedPajama dataset.
Training recipe.
To adapt $\mathsf{Inheritune}$ for this new setting, we perform step 1 and step 2 in Algorithm 1 without growing the model (i.e., we skip step 3). We use the first $n=13$ layers from our $k=26$ layer reference model. We call our small base LM Ours-1.5B(#tokens). We train our model with data repetition for eight epochs (each epoch uses all the 1B tokens) with a batch size of 131K tokens per batch. We use 1 A6000 GPU for less than half a day of training. The choice of training epochs is based on the analysis provided later in this paper (refer to Figure 13). We use the lit-gpt framework for training all small base LMs discussed in this paper. Further discussions on the training hyper-parameters can be found in the next Section.
Baseline models and evaluation.
We choose similarly sized (1-2B parameter) small base LMs trained with the RedPajama dataset and the reference base LM as primary baselines, as the quality of the pre-training data plays a key role in model development. We also include models OPT-1.3B Zhang et al. (2022) and Pythia-1.3B Biderman et al. (2023) as these models are pre-trained with a dataset similar to the RedPajama dataset. Table 6 lists the baseline models with their pre-training data.
In this study, we use few-shot accuracy, particularly 0-shot and 5-shot accuracy, on ten different downstream tasks to measure the quality of our 1.5B base LM. This evaluation of pre-trained base LLMs has been done in several prior works. Our evaluation methodology categorizes downstream tasks across four distinct genres: commonsense reasoning, natural language understanding, factuality, and natural language inference. We perform 0-shot evaluation for PIQA Bisk et al. (2020), BOOLQ Clark et al. (2019), WINOGRANDE Sakaguchi et al. (2020), WINOGRAD Kocijan et al. (2020), LOGIQA Liu et al. (2020), TruthfulQA Lin et al. (2022), MNLI Bowman et al. (2015), QNLI Wang et al. (2018) and WNLI Wang et al. (2018) datasets. Next, we perform a 5-shot evaluation on the massive multitask language understanding benchmark (MMLU) Hendrycks et al. (2020). We use the lm eval harness framework Gao et al. (2024) for the entire evaluation.
Model Commonsense Reasoning
Name (# train tokens) Reference Winograd PIQA Boolq WinoGrande Logiqa
OpenLLaMA-3B (1T) n/a 63.46 74.97 67.18 62.27 28.4
OPT-1.3B (300B) n/a 38.46 71.82 57.83 59.51 27.04
Pythia-1.4B (300B) n/a 36.54 70.89 63.12 56.99 27.65
MPT-1.3B (200B) n/a 63.46 71.44 50.89 58.09 28.26
Sheared LLaMA-1.3B (50B) LLaMA2-7B 36.54 73.45 62.02 58.17 27.34
Ours-1.5B (1B) OpenLLaMA-3B 50.96 56.47 61.68 51.69 25.19
Model Lang. Understanding & Inference Factuality
Name ( # train tokens) Reference MMLU(5) WNLI QNLI MNLI TruthfulQA
OpenLLaMA-3B (1T) n/a 27.21 50.7 51.3 37.3 35
OPT-1.3B (300B) n/a 24.96 42.25 51.29 35.82 38.67
Pythia-1.4B (300B) n/a 25.56 53.52 49.48 32.76 38.66
MPT-1.3B (200B) n/a 25.82 40.85 50.52 35.93 38.68
Sheared LLaMA-1.3B (50B) LLaMA2-7B 25.71 49.3 50.98 37.94 37.14
Ours-1.5B (1B) OpenLLaMA-3B 25.67 43.66 49.41 34.42 48.61
Table 5: Our 1.5B model achieves performance comparable to baseline models despite being trained with fewer tokens. Comparison of our target model ($\mathcal{M}_{\text{tgt}}$) derived using $\mathsf{Inheritune}$ with the reference model ($\mathcal{M}_{\text{ref}}$) and other baseline models of similar size when pre-trained from scratch and pre-trained with inherited weights and pruning. Although trained with fewer tokens, our model achieves performance comparable to the baseline models. We have highlighted all the scores in bold where our 1.5B model achieves at least 90% of the score compared to the reference LM or outperforms at least two of the baseline similar-size LMs. All the tasks are evaluated using 0-shot except MMLU, which is 5-shot. The models against which n/a is mentioned are trained from scratch.
B.1 Main Results in Low Data Regime
Table 5 presents a detailed performance evaluation across various tasks. Our 1.5B model, developed using $\mathsf{Inheritune}$, excels in 7 out of 10 individual tasks. It achieves a score of 90% or higher compared to the reference language model, which is twice its size and trained with 1000 times more data, or it outperforms at least two other base LMs of similar size trained with 50-300 times more data. Favorable scores are highlighted in bold.
Next, we compare our small LM with the MPT-1.3B [^4] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b model trained from scratch with 200B tokens of RedPajama dataset and find that we match 97% accuracy in all nine downstream tasks and the MMLU (5-shot) score. Additionally, we compare with OPT-1.3B and Pythia-1.3B models, showing that we outperform both in the MMLU (5-shot) score and perform comparably on the other nine datasets. This comparison illustrates that having a large reference base LM and a subset of its pre-training data allows the inherited target size base LM to be trained remarkably more sample-efficiently than training from scratch. Extended discussions on comparisons with the ShearedLLaMa model, generated by pruning and continual training from LLaMA2-7B, are provided in the supplementary materials.
Ablation of Inheritune Across Different Model Sizes with 1B Tokens.
In the previous section, we considered a single choice of $n=k/2$, i.e., half the layers, for the size of the smaller model. Here, we investigate $\mathsf{Inheritune}$ with different choices of $n$, but the same 1B token dataset). All models use OpenLLAMA-3B as the large pre-trained reference model, with consistent training hyperparameters, changing only the choice of $n$.
We developed eight different submodels with $n=\{4,6,8,10,13,16,18,20\}$. Figure 11 shows the MMLU (5-shot) score as a function of $n$. As expected, the trend line is positive-sloping. The submodel with 20 layers slightly decreases performance, potentially due to data overfitting as the model size increases. The training details for all these submodels are consistent with the target 1.5B small base LM and are detailed in the appendix. A more comprehensive investigation on the choice of $n$—including varying both $n$ and the number of training tokens jointly and evaluating a broader set of tasks is left for future work.
Figure 11: Inheritune scales across multiple different model sizes. Utilizing the OpenLLaMA-3B as a reference large base LM, demonstrates that multiple performant small base LMs of target size can be crafted using $\mathsf{Inheritune}$ with just 1B training tokens. The MMLU (5-shot) as a function of the number of submodels.
B.2 Additional analysis with larger reference LMs and 50B data
We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available. Initially for the main results we limited ourselves to 1B (i.e. 0.1%) tokens from the 1T pre-training data, here we use a 50B subset (i.e. 5%) of the pre-train data. Moreover we also extend this study to include larger base LMs of 7B parameters as reference models, employing OpenLLaMA-7B and LLaMA2-7B as reference models. For the purpose of this study we do not repeat the tokens from our 50B subset. As shown in Figure 12, we observe that there is clear improvement in overall MMLU (5-shot) score with more data. Additionally it is interesting to see that 1.5B (or 1.6B models) developed with $\mathsf{Inheritune}$ using larger reference models show even greater improvements when fed with 50B subset of non repetitive data (i.e fresh tokens). We present a Table 8 using Figure 12 to show the best MMLU (5-shot) scores achieved using different reference LMs. For developing our small base LMs using larger reference LMs we use $n$=7 (i.e. 7 layers). The training details are discussed in the following section.
Model Training Data (# tokens)
OpenLLaMA-3B v1(ref) RedPajama(1T)
Ours-1.5B* RedPajama (1B)
Shear-LLaMA-1.3B* RedPajama(50B)
MPT-1.3B RedPajama(200B)
Pythia-1.4B The Pile(300B)
OPT-1.3B Custom data(300B)
Table 6: Comparison of training data across baseline models. Overview of reference and baseline models, including their pre-training datasets and the number of tokens used during training. Note the significant variation in training data size, ranging from 1B to 1T tokens.
Models (# train tokens) GPU Count GPU Type Time (# days)
MPT-1.3B (200B) 440 A100 half
Pythia-1.4B (300B) 64 A100 4.6
TinyLLaMA-1.1B (3T) 16 A100 90
OPT-1.3B (300B) 992 A100
Sheared LLaMA-1.3B (50B) 16 A100
OpenLLaMA-3B (1T) 256 TPU v4 10
Our-1.5B (1B) 1 A6000 ~half
Table 7: Computational efficiency of Inheritune versus baseline models. Comparison of pre-training compute requirements for publicly available small base LMs and our $\mathsf{Inheritune}$-derived model. Metrics include GPU count, GPU type, and training duration, highlighting Inheritune’s significant reduction in computational resources.
Figure 12: Impact of reference model choice on Inheritune performance. MMLU (5-shot) scores for 1.5B base LMs derived using $\mathsf{Inheritune}$, trained on 50B unique tokens. Comparison across three reference models: OpenLLaMA-7B, LLaMA2-7B, and OpenLLaMA-3B. Results demonstrate Inheritune’s effectiveness with various large language models as references.
Figure 13: Performance of our 1.5B base LM derived using $\mathsf{Inheritune}$ based on existing OpenLLaMA-3B base model. Here we use 1B tokens and perform data repetition (epochs) during training. We further evaluate our model on an average of 9 different datasets (left) and MMLU benchmark (right).
Model (# tokens), ref MMLU(5)
Ours-1.6B (1B), LLaMA2-7B 24.27
Ours-1.5B (1B), OpenLLaMA-3B 25.67
Ours-1.5B (50B), OpenLLaMA-3B 25.71
Ours-1.6B (50B), LLaMA2-7B 26.07
Ours-1.6B (50B), OpenLLaMA-7B 26.72
Table 8: Performance comparison of models on the MMLU (5-shot) task. Our models, even when trained with fewer tokens, show competitive performance compared to benchmarks. We have highlighted the best MMLU 5-shot score in bold.
Model (# tokens) Data type MMLU (5-shot)
Ours-1.5B (1B) 10 epochs 24.95
Ours-1.5B (50B) 10B fresh 23.62
Ours-1.5B (1B) 20 epochs 25.46
Ours-1.5B (50B) 20B fresh 24.96
Table 9: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 12 and 13.
Ablations with number of epochs.
We ran ablations (refer Figure 13) to choose the total number of epochs (multiple passes over the data) and observe that repetition when training our 1.5B (or 1.6B) LM is helpful particularly for MMLU. We also observe that the for an average of all the 9 other datasets (i.e. except MMLU) peaks it’s performance at 5 epochs and then deteriorates. Some prior works have studied this phenomenon that the scaling of downstream tasks with data is not always linear Biderman et al. (2023).
To repeat or not to repeat the tokens.
Next we tackle the question – whether one should re-use 1B tokens for multiple epochs or use the same number of fresh tokens? Some prior works have recommended that if you have a reasonably large size dataset one can repeat it upto 4 epochs Muennighoff et al. (2023). In our study we observe that one can safely re-use 1B tokens upto 10-20 epochs as shown in Table 9. We emphasis that this phenomenon needs a through investigation in itself and we defer this to future work. The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
B.3 Implications of Low Data Regime
In this section, we discuss some of the key implications of our work in low data regime.
Cheap and easy development of small base LMs.
Pre-training a small base LM of 1-2B parameters from scratch is extremely expensive. For instance mpt-1.3B base LM is pre-trained with 440 A100 GPUs for half a day, while the Pythia-1.4B base LM Biderman et al. (2023) utilized 64 A100-40GB GPUs for 4.6 days. Similarly, TinyLLaMA-1.1B model Peiyuan Zhang & Lu (2023) was pre-trained using 16 A100 GPUs for 3 months. Our 1.5B (1B data variant) LM shows competitive performance despite being trained with 1 A6000 GPU for less than 12 hours. The computational details are provided in Table 7, comparing the training resources of the baseline models listed in this paper. Typically small base LMs are finetuned for a specific task before deployment and are not used in it’s base form. With $\mathsf{Inheritune}$ we present a really easy and cheap way for developing a small base LM to be later finetuned before deployment.
Naive baseline for pre-training a scaled down variant of large base LMs.
Typically small variants of large base LMs are pre-trained using the same pre-training data Peiyuan Zhang & Lu (2023); Groeneveld et al. (2024). Our recipe introduces a new perspective of identifying sufficient depth without losing any generalization on the held out validation set. Next, we also show that even with a small fraction of pre-train data (randomly sampled) and few initial layers of the large base LM one can develop a small base LM. Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
Appendix C Implementation Details
C.1 Training details of GPT-2 models
For our main experiments, we focused on three sizes of GPT-2 models Radford et al. (2019): the vanilla GPT-2 xlarge with 1.5B parameters, GPT-2 large with 770M parameters and the vanilla GPT-2 medium with 355M parameters. We developed several variants of these models by adjusting the number of layers and hidden size. We trained all GPT-2 models with data repetition while using OpenWebText dataset, the trainset has 9B tokens and the validation set has 4.4M tokens. The key architectural configurations for the reference models, our models, and baseline models discussed in this paper are summarized in Table 10.
For all training runs, we used GELU activations, disabled bias terms, and removed dropout, following the nanoGPT codebase and Liu et al. (2023). We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$. The GPT-2 models were trained on a single node with 3 A100 GPUs (each with 40 GB of memory) using distributed data parallelism and gradient accumulation. In line with Liu et al. (2023), we scaled the attention logits inversely to the layer index across all GPT-2 models. Most hyperparameters were adapted from Liu et al. (2023), with key details provided below.
Hyper-parameter details of GPT-2 Medium and variants.
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 large and variants.
  • Batch size: 16K tokens
  • Learning rate: $2\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 xlarge and variants.
  • Batch size: 16K tokens
  • Learning rate: $1.5\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of knowledge distillation training.
We use the below loss for as our distillation based training loss. The validation loss is the student_loss.
[…]
  • Model: 16-layer and 12-layer GPT-2 medium variants
  • Softmax temperature: 1
  • $\alpha$: 0.6
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 50K
Models Layers Hidden Size Heads Variant
[…]
GPT2-large(770M) 36 1280 20 Original
GPT2-large(680M) 32 1280 20 Original
GPT2-medium(355M) 24 1024 16 Original
[…]
GPT2-medium 16 512 8 half width
[…]
GPT2-large 18 1280 20 Ours
GPT2-large 16 1280 20 Ours
GPT2-medium 16 1024 16 Ours
Table 10: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
C.2 Training details of 1.5B OpenLLaMA model
Small base LMs trained with 1B data
We present our main results with Our-1.5B model trained with an existing OpenLLaMA version 1 Geng & Liu (2023) and 1 B tokens randomly sampled from 1T redpajama version1 data. The hyper-parameters related to this model is provided below. It is important to note that our claim that we only use 1 GPU for less than 12 hours to train Our-1.5 B model is specific to models derived using $\mathsf{Inheritune}$ with 1B data. Next we also train multiple sub-models as shown in Figure 11 the training details remains consistent with that of the initial model discussed earlier. However we observe that increasing the number of layers in a sub-model also increase the training time.
Hyper-parameter details of our 1.5B base LM derived using OpenLLaMA-3B as refernce LM:
  • Training tokens: 1B
  • Training epochs: 8
  • Training steps: 64K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm up steps: 1000
  • Batch size: 131K
  • GPU count: 1
  • GPU type: A6000
  • GPU hours: $\sim$ 8 hours
  • GPU hours/epoch: $\sim$ 54 minutes
Training details of small base LMs with 50B data.
We also trained our 1.5B model with larger subsets of data as shown in Figure 12.It is important to note that all the intermediate tokens until 50B are intermediate checkpoints of a single training run. Some of the key hyper-parameters of our training runs are discussed below. We have also trained three variants of small base LMs utilizing 3 different reference base LMs namely OpenLLaMA-3B, OpenLLaMA-7B and LLaMA2-7B. For target LMs developed with OpenLLaMA-3B we use $n$=13 i.e. 13 layers. For target LMs developed using reference LMs of 7B parameters we use $n$=7 i.e. 7 layers. The training hyper-parameters remains consistent across all the models trained with 50B subset of the pre-train data.
Training hyper-parameters of our target 1.5B and 1.6B small base LMs:
  • Training tokens: 50B
  • Training epochs: $\sim$1
  • Training steps: 191K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm-up steps: 1000
  • Batch size: 131K tokens
  • GPU count: 1
  • GPU type: A100
  • GPU hours: $\sim$18 hours