### v3

# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

Abstract
Large Language Models (LLMs) leverage the transformer architecture and its self-attention mechanism to achieve remarkable performance. However, we identify a structural inefficiency in standard pre-trained decoder-style LLMs: attention matrices in many deeper layers frequently degenerate, often collapsing to near rank-one, single-column structures. These layers, which we term lazy layers, are redundant and inefficient.
To address this, we introduce $\mathsf{Inheritune}$, a simple yet effective training recipe for developing smaller, more attentive, and high-performing language models. $\mathsf{Inheritune}$ initializes a compact model by inheriting the potent early layers from a larger pre-trained model, then progressively retrains and expands it. Our demonstration of $\mathsf{Inheritune}$’s efficacy on various models and datasets reveals that models trained with this approach, despite having significantly fewer layers, can achieve performance comparable to or even exceeding their larger counterparts. This approach enables model compression by design, producing language models that are both compact and high-performing. Code is available at https://github.com/sanyalsunny111/LLM-Inheritune.
1 Introduction
Large Language Models (LLMs) are built with decoder-style transformer blocks (Vaswani et al., 2017). These models are typically designed to be large, with a significant portion of their parameters dedicated to their depth, with multiple transformer blocks stacked with eachother building model capacity. Each block or layer in the stack refines the representations learned by the previous blocks, allowing the model to develop a nuanced understanding of the input data. As these models scale in depth and size, their performance tends to improve Kaplan et al. (2020); Hoffmann et al. (2022), benefiting from increased model capacity.
The causal self-attention (hereafter referred to as attention) mechanism is arguably the most crucial component of a transformer block. It allows models to combine tokens as a weighted linear sum of their attention scores, effectively capturing long-range dependencies and contextual relationships within text data. However, as models grow in depth, they often encounter a phenomenon known as attention degeneration caused by collapse in the attention rank ((Noci et al., 2022; Dong et al., 2021; He et al., 2023)). Notably, this phenomenon has not been studied in the context of standard LLMs. A formal discussion on attention degeneration is provided in Section 2.
In this paper, we empirically analyze 24-layer GPT-2 medium and 36-layer GPT-2 large models (decoder-style LLMs) Radford et al. (2019) for attention degeneration and observe that many deeper layers in both models exhibit rank-1 attention matrices. Further investigation reveals that most of these rank-1 matrices are also single-column, i.e. their mass is concentrated to a single column. Our attention matrix analysis is presented in Figure 1. We term these deeper layers, where all attention matrices of a given layer are degenerated, as lazy layers.
Motivated by this new finding we aim to develop performant small base language models (LMs) utilizing weights from inefficient larger base LMs. A base LM is a decoder-style model trained solely for next-token prediction without additional enhancements like instruction tuning or reinforcement learning with human feedback (RLHF). Our proposal is straightforward, we start by initializing our smaller LM (target) using the first few blocks from a large pre-trained LM (reference). We then train the target model for a specified number of steps. After this initial training, we incrementally grow the target model by adding more blocks, continuing the training process until it matches or surpasses the pre-train validation loss (also val loss) of the reference model. During the growth phase, the newly added blocks can be initialized with lazy layers of the reference LM. We refer to this simple yet effective training approach as $\mathsf{Inheritune}$.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 XLarge (1.5B) L=48, H=1600
Figure 1: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near rank-1, limiting their ability to learn meaningful representations. We compute $\text{MaxRank}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set. Our rank analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models reveals that attention matrices in many deeper layers collapse to near rank-1.
In summary, our key contributions are as follows:
  1. 1.
    Analysis of Attention Degeneration Leading to Lazy Layers. We empirically investigate attention degeneration in standard LLM settings. Our analysis shows that rank-collapsed attention matrices often exhibit single-column structures, revealing a significant structural inefficiency in the attention mechanism of standard LLMs in deeper layers (refer Figure 1).
  2. 2.
    Introduction of Inheritune. We propose $\mathsf{Inheritune}$ as an approach to effectively train high-performing, smaller models. This method involves inheriting a few early blocks from a larger pre-trained model and progressively growing and training the smaller model. The initialization is entirely zero-shot. We validate the effectiveness of $\mathsf{Inheritune}$ through comprehensive experiments using GPT-2 xlarge (1.5B), GPT-2 large (770M), and GPT-2 medium (355M) models, trained on the OpenWebText and the FineWeb_edu dataset.
  3. 3.
    Evaluation Against Multiple Baselines. Models derived using $\mathsf{Inheritune}$ consistently outperform various baselines, including much larger models trained from scratch (refer Figure 5 and Figure 6), model initialization and efficient training baselines (refer Table 1), and models trained using two knowledge distillation techniques (refer Figure 7).
2 Attention Collapse and the Emergence of Lazy Layers in LLMs
Preliminaries:
A vanilla transformer-based model consists of $L$ transformer blocks (layers). The model operates on an input sequence $X\in\mathbb{R}^{T\times e}$, where $T$ denotes the sequence length (number of tokens), and $e$ represents the embedding dimension or model hidden size. The output of each layer $l$ is denoted as $X^{(l)}\in\mathbb{R}^{T\times e}$.
Each transformer block primarily consists of two sub-layers: a self-attention block and a position-wise feed-forward network (FFN). The self-attention mechanism enables the model to weight the relevance of different tokens in the sequence relative to each other. Specifically, for a single attention head, the attention computation is defined as $\text{Attention}(Q,K,V)=\underbrace{\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)}_{\textbf{Attention matrix: $A(X)$}}V$
where the queries $Q=XW_{Q}$,  keys $K=XW_{K}$,  and values $V=XW_{V}$ are linear transformations of the input $X$. Here, $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{k}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ are the weight matrices for the queries, keys, and values, respectively. Typically, $d_{k}=d_{v}=\frac{d}{h}$, where $h$ is the number of attention heads. In this single-head scenario, we set $d_{k}=d_{v}=d$.
The attention matrix $A(X)\in\mathbb{R}^{T\times T}$ captures the pairwise attention scores between all token positions in the sequence. The softmax is applied row-wise. The attention matrix $A(X)$ is then used to compute a weighted sum of the value vectors. Attention rank collapse refers to the phenomenon where the attention matrices $A(X)$ of individual heads in many layers of transformer-based language models lose their expressive capacity, converging towards lower effective rank structures. Specifically, the effective rank of attention matrices significantly reduces, often approaching rank-1, limiting the model’s ability to meaningfully differentiate between token interactions across positions in the sequence. Previous research by Dong et al. (2021) and He et al. (2023) has shown that in self-attention networks (SANs) without residual connections and feed-forward networks (FFNs), the rank of an attention matrix converges to rank-1 doubly exponentially with respect to the depth of the model. This phenomenon, known as rank collapse of attention matrices, results in a loss of expressive power as the attention mechanism attends to all tokens uniformly. Noci et al. (2022) showed that even with residual connections (without layernorm) attention matrices can still lose rank in deeper layers if the residual connections are not scaled by $1/\sqrt{L}$. Interestingly they also linked the rank collapse to vanishing gradients of the keys and queries in deeper layers which affects the overall trainability of the transformer based models. However, these findings do not directly apply to the standard LLMs, as transformer blocks in these models include residual connections, layernorms and FFNs, which are expected to mitigate both rank collapse and the vanishing gradient problem.
Approximate Rank Computation of Attention Matrices
To assess the presence and severity of rank collapse within standard decoder style transformer architectures (e.g., GPT-2, LLaMA etc.), we utilize singular value decomposition (SVD) for each attention matrix $A(X)=U\Sigma V^{\top},$ where $\Sigma$ is a diagonal matrix containing singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{T}\geq 0$. The approximate rank (referred to as rank hereafter) of an attention matrix, parameterized by a variance threshold $\tau$, is formally computed as:
$k^{*}=\min\left\{k\in\{1,2,\dots,T\}\mid\frac{\sum_{i=1}^{k}\sigma_{i}^{2}}{\sum_{j=1}^{T}\sigma_{j}^{2}}\geq\tau\right\},$
where $\tau\in(0,1)$ represents the proportion of variance that must be captured by the top $k$ singular values. A lower value of $k^{*}$ indicates stronger rank collapse. In this work, we set $\tau=0.90$. A lower $k^{*}$ indicates stronger rank collapse.
In Figure 1, we present the layer-wise analysis of rank of GPT-2 models. For this analysis, we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText with 4M tokens, each with a sequence length of $T=100$ tokens across all attention heads within each layer. We then define the average approximate rank for each head and layer as $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$. Subsequently, we aggregate this metric per layer by taking the maximum rank across heads: $\mathrm{MaxRank}^{(l)}=\max_{h}\{\mathrm{Rank}^{(h,l)}\}$. As demonstrated in Figure 1, $\text{MaxRank}^{(l)}$ reveals that many deeper layers exhibit attention matrices that are predominantly near rank-1. We highlight that this rank collapse occurs in GPT-2 medium, Large, and xLarge models—which are standard architectures incorporating skip connections and FFNs—thereby extending the findings of Dong et al. (2021) and Noci et al. (2022) to these widely-used LLMs. The rank analysis is quite robust to the dataset in Figure 10, we performed the same analysis using the same models on held out data from Fineweb_edu dataset (Penedo et al., 2024). Overall, the degeneration of attention matrices in deeper layers provides quantitative evidence for the existence of lazy layers. Specifically, we observe that some deeper layers exhibit a near-complete rank collapse of attention matrices across all heads, suggesting potentially reduced representational capacity and less effective token mixing in these layers.
2.1 The Functional Ineffectiveness of Lazy Layers
Having identified lazy layers, we investigate their practical utility: Do these structurally degenerated layers retain transferable knowledge, or are they functionally impaired? Our experiments suggest the latter.
(a) Rank profile (GPT-2 Small 12L).
(b) Performance of 4L variants.
Figure 2: Higher-rank layers transfer better. (Left, 2(a)) Layer-wise $\mathrm{MaxRank}^{(l)}$ of a pre-trained 12L GPT-2 Small. (Right, 2(b)) Validation loss of 4L variants initialized with potent blocks (AvgRank $\approx 8.4-9.5$) vs. a lazy block (AvgRank $\approx 1.2$) or random weights, after 100K steps. Lazy block initialization mirrors random.
In the first set of experiments, we trained a vanilla GPT2-small (125M) model with 12 layers for 100K steps on the OpenWebText dataset. We then performed the rank analysis described earlier, with results presented in Figure 2. Specifically, we aggregated the approximate ranks over groups of contiguous layers using $\text{AvgRank}=\frac{1}{L}\sum_{l=1}^{L}\text{MaxRank}^{(l)}$, where $L$ is the number of layers in each group. Subsequently, we trained three GPT2-small variants [^1] A variant shares the same configurations as the parent model but has fewer layers. for 100K steps, each initialized with a different contiguous block of four layers from the trained vanilla GPT2-small model: (a) layers 1–4, with $\text{AvgRank}=8.40$; (b) layers 5–8, with $\text{AvgRank}=9.48$; and (c) layers 9–12, with $\text{AvgRank}=1.22$. The last model is initialized with lazy layers. For comparison, we also trained another GPT2-small variant with random initialization for 100K steps. All models were trained on the OpenWebText dataset. As shown in Figure 2, the model initialized with layers from the vanilla GPT2-small model having higher AvgRank demonstrated the best performance (i.e., lowest final validation loss). Additionally, we observed that the model initialized with lazy layers performed very similarly to the model with random initialization suggesting that lazy layers contain minimal transferable knowledge. The key results are also summarized in Table 4.
(a) GPT2-medium (12-layer) variants.
(b) GPT2-large (16-layer) variants.
Figure 3: When initializing 12-layer and 16-layer variants of GPT2-medium and GPT2-large with deeper (lazy) layers showing degenerated attention, performance is comparable to random initialization. In contrast, early-layer initialization leads to significantly better convergence and generalization.
For the second set of experiments we utilized larger models namely GPT-2 medium and GPT-2 large both similarly trained for 100K steps using OpenWebText dataset. Here we initialized a 12-layer GPT-2 medium variant and an 18-layer variant of GPT-2 large using lazy layers extracted from pre-trained 24-layer GPT-2 medium and 36-layer GPT-2 large models. We then trained these GPT-2 variants on the same dataset for 10K steps. For comparison, we conducted two baseline experiments where the GPT-2 variants were initialized either with the first half of transformer layers (potent layers with high AvgRank) and with random initialization. As shown in Figure 3, models initialized with lazy layers demonstrate poor transferability, performing similarly to models with random initialization. This provide additional evidence that lazy layers with fully degenerated attention, fails to learn meaningful representations.
3 Inheritune: Our Proposed Training Recipe
This section provides a detailed description of our method, key implementation considerations, and how it addresses the inefficiencies present in current architectures.
  1. 1.
    Inherit: Initialize $\mathcal{M}_{\text{tgt}}$ with the first $l$ out of $L$ layers of $\mathcal{M}_{\text{ref}}$, including prediction head, and token embeddings.
  2. 2.
    Train: Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps on $\mathcal{D}_{\text{train}}$ and evaluate on $\mathcal{D}_{\text{val}}$.
  3. 3.
    Grow: If needed, increase $\mathcal{M}_{\text{tgt}}$’s size by adding next few contiguous layers and repeat steps 1-2 until desired performance is achieved.
As previously established, we have identified the problem of attention degeneration and its connection to lazy layers, highlighting specific inefficiencies in pre-trained LLMs. In this work, we transform this challenge into an opportunity to create smaller base language models, which we refer to as target models $\mathcal{M}_{\text{tgt}}$, that achieve comparable performance with similar or lower validation loss compared to their larger, less efficient counterparts, which we term reference models $\mathcal{M}_{\text{ref}}$.
Algorithm 1 Inheritune: Training Recipe for Small Language Models
1:Reference model $\mathcal{M}_{\text{ref}}$ with $L$ layers, datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$, steps $\mathsf{T}$
2:Copy embedding layer and LM head from $\mathcal{M}_{\text{ref}}$ to $\mathcal{M}_{\text{tgt}}$
3:Select $l$ early contiguous layers from $\mathcal{M}_{\text{ref}}$ with high AvgRank
4:Initialize $\mathcal{M}_{\text{tgt}}$ with selected layers between embeddings and LM head
5:Train $\mathcal{M}_{\text{tgt}}$ on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps
6:while $\mathcal{M}_{\text{tgt}}$ performance < $\mathcal{M}_{\text{ref}}$ performance on $\mathcal{D}_{\text{val}}$ do
7:     Grow $\mathcal{M}_{\text{tgt}}$ by inheriting additional layers
8:     Train $\mathcal{M}_{\text{tgt}}$ for $\mathsf{T}$ steps
9:end while
10:return Optimized model $\mathcal{M}_{\text{tgt}}$
Our proposed solution builds on two key insights: (1) the early layers of deep LLMs contain a higher concentration of potent layers with high AvgRank values, making them suitable for effective model initialization, and (2) lazy layers can be identified, removed, or utilized in smaller numbers, then subsequently re-trained to improve overall model capacity.
Figure 4: Overview of the $\mathsf{Inheritune}$ training recipe using a 24-Layer GPT-2 medium model example. A smaller target model is initialized using early layers from a larger, pre-trained reference model. The target model goes multiple rounds of training while inheriting more early layers until it matches the reference model. The intensity of the red color in layers correlates with $\mathrm{MaxRank}^{(l)}$.
Setup:
We split the dataset into a training set $\mathcal{D}_{\text{train}}$ and a validation subset $\mathcal{D}_{\text{val}}$. Next, we assume that there exists a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $L$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{L-1}\}$ trained with $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$ steps. We want to train a smaller model $\mathcal{M}_{\text{tgt}}$ with the same or better validation loss (lower is better) compared to its larger counterpart $\mathcal{M}_{\text{ref}}$.
We now present $\mathsf{Inheritune}$, our proposed training recipe for efficiently developing small base language models (LMs). $\mathsf{Inheritune}$ operates on the principle of zero-shot initialization and progressive growth. The $\mathsf{Inheritune}$ process consists of three main steps, which we present below and formalize in Algorithm 1:
With our method now formally described, we turn to empirical validation. In the following sections, we present comprehensive results demonstrating $\mathsf{Inheritune}$’s effectiveness across various scenarios, including different model sizes and data regimes. In addition, we conducted an in-depth ablation study to analyze the impact of initialization on performance, providing insights into the adaptability of our approach.
4 Experiments
We evaluate $\mathsf{Inheritune}$ through a comprehensive set of experiments using several GPT-2 models: 48-layer GPT-2 xlarge (1.5B), 36-layer GPT-2 large (770M), 32-layer GPT-2 large (668M) and 24-layer GPT-2 medium (355M) models Radford et al. (2019). Table 11 provides detailed specifications of all model configurations used.
We use two training datasets in our experiments: OpenWebText Gokaslan & Cohen (2019) and Fineweb_edu Penedo et al. (2024). Our experimental setup closely follows prior works Liu et al. (2023); Sanyal et al. (2024). For models trained on OpenWebText, we report validation loss (log perplexity) on a held-out set. For those trained on Fineweb-Edu, we report training loss (also log perplexity). In addition, we conduct zero-shot downstream evaluations using the lm-evaluation-harness Gao et al. (2024) across five standard tasks: ARC-easy (ARCE; Clark et al. (2018)), LAMBADA Paperno et al. (2016), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020). We also perform a thorough ablation study of our initialization strategy, focusing on 16-layer GPT-2 medium variant.
We provide experimental details our $\mathsf{Inheritune}$ training recipe using a GPT-2 medium model as an example; similar procedure was applied to train other models. A visualization of the training recipe is presented in Figure 4. Our recipe for applying $\mathsf{Inheritune}$ involves the following steps.
  1. 1.
    Reference Model: We train the full 24-layer GPT-2 medium model on $\mathcal{D}_{\text{train}}$ for 100K steps and evaluate its validation loss ( log-perplexity) on $\mathcal{D}_{\text{val}}$. This establishes our benchmark validation loss.
  2. 2.
    Model initialization: We initialize an 12-layer model ($l=L/2$) using the reference model.
  3. 3.
    Training and Evaluation: We train the 12-layer model on $\mathcal{D}_{\text{train}}$ for $\mathsf{T}$=100K steps and evaluate its validation loss.
  4. 4.
    Iterative Refinement: If the smaller model’s performance is inferior to the reference model, then we incrementally increase its size by two layers and repeat steps 2-3 until we achieve parity with the reference model’s validation loss.
We choose $l=L/2$ as the starting point and increase the model size by two layers in each round across all our experiments, aiming to minimize the number of training rounds. In principle, $\mathsf{Inheritune}$ should generalize to other hyperparameter choices as well.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 5: Models derived using Inheritune converge faster and match the final validation loss of the full-sized model, despite having much fewer layers. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge variant, 18-layer GPT-2 Large variant, 16-layer GPT-2 Medium variant) against their full-sized counterparts and same sized variants trained from scratch. All models are trained for 100K steps using OpenWebText data with data repetition.
(a) GPT2-Medium
(b) GPT2-Large
Figure 6: Models derived using $\mathsf{Inheritune}$ converge faster and match the final validation loss and average downstream performance (Perf.) across 5 datasets of the full-sized model, while outperforming same-sized models. (Plots) Validation loss curves for our 16-layer variants versus full-sized models. (Table) Zero-shot downstream performance comparison. All models are trained for 100K steps using FineWeb_edu data without data repetition.
Models Recipe Layers Perf. Medium rand init 24 47.74 rand init 16 46.84 Ours 16 48.08 Large rand init 32 49.44 rand init 16 48.17 Ours 16 49.75
Baselines.
We compare GPT-2 model variants (variants are models with lesser layers compared to their vanilla configuration) trained using $\mathsf{Inheritune}$ against two types of baselines: (1) larger reference models with more layers, and (2) similar-sized models with the same number of layers trained from scratch (randomly initialized). Additionally, we compare $\mathsf{Inheritune}$ against efficient training baselines that leverage pre-trained model initialization. These include stacking Gong et al. (2019); J. Reddi et al. (2023), hybrid stacking (which initializes the model using layers from the reference model), and half-width initialization (where all layers are retained, but the hidden dimension and number of attention heads are halved and initialized from the reference model). Finally we briefly compare $\mathsf{Inheritune}$ with knowledge distillation (Hinton et al., 2015). Detailed descriptions of these baselines are provided in the supplementary material Section B.
Models Layers Recipe Steps Pre-train Val loss ($\downarrow$)
GPT-2 Medium 24 half-width 100K 3.04
16 stacking 100K 2.84
16 hybrid-stacking 100K 2.83
16 Ours 100K 2.81
GPT-2 Large 36 half-width 100K 3.06
18 stacking 100K 2.87
18 hybrid-stacking 100K 2.89
18 Ours 100K 2.80
GPT-2 xLarge 48 half-width 100K 2.77
24 stacking 100K 2.65
24 hybrid-stacking 100K 2.64
24 Ours 100K 2.64
Table 1: $\mathsf{Inheritune}$ outperforms baseline zero-shot initialization and efficient training techniques. Comparison of pre-training validation loss for GPT-2 xLarge, GPT-2 Large and GPT-2 Medium variants. $\mathsf{Inheritune}$-derived models consistently achieve lower loss compared to models initialized with stacking, hybrid stacking, and half-width techniques.
4.1 Results and Discussions
Models trained with $\mathsf{Inheritune}$ outperform much larger models trained from scratch.
We present our main results in Figure 5. The 24-layer, 18-layer, and 16-layer variants derived using $\mathsf{Inheritune}$ from the vanilla 48-layer GPT-2 xlarge, 36-layer GPT-2 large, and 24-layer GPT-2 medium, respectively, achieve comparable or lower validation losses than both their full-sized counterparts and same-sized models trained from scratch, when trained for the same number of steps (100K). Our GPT-2 xlarge and GPT-2 large variants require a single round of $\mathsf{Inheritune}$ training, while the GPT-2 medium variant undergoes three rounds with 12-, 14-, and 16-layer configurations. From a convergence perspective, prior work has linked overparameterization to faster convergence Bengio et al. (2005); Vaswani et al. (2018). Interestingly, we find that smaller models derived using $\mathsf{Inheritune}$ converge just as fast as their larger counterparts. A detailed summary of these results is provided in the appendix Table 3.
We conducted additional training experiments without data repetition, following standard LLM pre-training practices as discussed in Touvron et al. (2023); Biderman et al. (2023) using the high-quality Fineweb_edu dataset. We trained a 32-layer GPT-2 large (668M) and a 24-layer GPT-2 medium (355M) reference model from scratch. Next, we trained two 16-layer variants: one derived from GPT-2 large and the other from GPT-2 medium, using their respective reference models following $\mathsf{Inheritune}$. For comparison, we also trained 16-layer baseline models from scratch. All models were trained for 100K steps. We observe thematically consistent results: as shown in Figure 6, the 16-layer variants trained with $\mathsf{Inheritune}$ consistently match the performance of their full-sized counterparts and outperform same-sized baselines, both in terms of training loss and zero-shot downstream evaluation. Downstream results are averaged over five tasks (see table within Figure 6); the complete breakdown is provided in Table 5 placed in appendix. Model configurations and training hyper-parameters are detailed in the supplementary material (refer Section F).
Figure 7: A 16-layer GPT-2 medium variant derived using $\mathsf{Inheritune}$ converges faster and generalizes better than a same-sized model trained with logit-based distillation baselines.
Models trained with Inheritune outperform all zero-shot model initialization baselines.
In Table 1, we compare GPT-2 xlarge, GPT-2 large, and GPT-2 medium variants trained with $\mathsf{Inheritune}$ against same-sized variants trained with stacking, hybrid, and half-width initialization baselines. The half-width baseline performs poorly, revealing the limitations of naive width reduction. While stacking and hybrid stacking demonstrate reasonable performance, they still fall short compared to $\mathsf{Inheritune}$. Across all cases, $\mathsf{Inheritune}$ consistently outperforms these baselines, highlighting its effectiveness as an initialization strategy. For a detailed view of the training dynamics across all methods, refer to the training curves in supplementary Figure 14.
Distillation vs Inheritune.
In Figure 7, we compare 16-layer GPT-2 medium variants derived using vanilla knowledge distillation Hinton et al. (2015) and DistillBERT-style distillation Sanh et al. (2019), which leverages teacher layers for model initialization, vanilla training from scratch and $\mathsf{Inheritune}$. Our model demonstrates faster convergence and significantly better final generalization after 50K training steps. Additional distillation experiments can be found in the supplementary materials.
4.2 Ablations
Layers Initialization Steps Val loss ($\downarrow$)
16 attn (w/ layernorm) 100K 2.84
16 mlp (w/ layernorm) 100K 2.85
16 attn+mlp (w/o layernorm) 100K 2.80
16 Ours (attn+mlp w/ layernorm) 100K 2.81
Table 2: Impact of initializing various sub-modules within a transformer block. We compare validation loss of a 16-layer GPT-2 medium variant when different sets of sub-modules are initialized with weights from the first 16 layers of a 24-layer GPT-2 medium reference model. All models are trained on the OpenWebText dataset. Key findings: (1) $\mathsf{Inheritune}$ initialization and attention + MLP initialization result in similar performance improvements; (2) layernorm initialization shows minimal impact. A detailed training curve is shown in Figure 15.
We conducted extensive experiments to better understand which sub-module initializations within a transformer block lead to improved generalization (in terms of validation loss) and faster convergence. For these ablations, we fixed the model to a 16-layer GPT-2 medium variant and explored three different sub-module initializations using weights from a 24-layer GPT-2 medium reference model. We initialize the transformer blocks with 1) attention ((key, query, value, and projection) and the layernorm [^2] In GPT-2 models layernorm blocks are parameterized. weights (attn w/ layernorm), 2) attention and mlp weights without the layer-norm (attn+mlp w/o layernorm), and 3) mlp weights with the layer norm (mlp w/ layernorm). We emphasize that $\mathsf{Inheritune}$ performs initialization by inheriting attention and mlp weights with the layer norm (attn+mlp w/ layernorm).
As shown in Table 2, models trained with attention and mlp weights demonstrated the best performance, regardless of the layer norm initialization. A detailed validation loss vs training steps plot is presented in supplementary Figure 15. We conclude that initializing both attention and MLP weights provides a clear advantage. Surprisingly, we also observed that initializing either the attention or mlp weights resulted in similar improvements in both convergence speed and final validation loss.
5 Related Works
Attention degeneration has been studied in the past through the lens of attention rank collapse Dong et al. (2021) leading to representation collapse, and attention entropy collapse Zhai et al. (2023) leading training instability. This also has been studied is a theoretical setup for transformer models by Noci et al. (2022); Barbero et al. (2024). Recently He et al. (2023) address rank collapse in self-attention networks (SANs) without residual connections or layer norms, using two model initialization techniques that enable faithful signal propagation—i.e., $\Sigma_{L}$ of $A(X^{L})$ does not collapse in deeper layers. However, this approach significantly slows down training. Noci et al. (2022) proposes scaling residual connections by $1/\sqrt{L}$, while Barbero et al. (2024) suggest that adding additional tokens to already long sequences of repeated tokens can help mitigate degeneration. In contrast to prior works, we address attention degeneration by developing smaller models that eliminate structural inefficiencies and training these models to match the performance of their larger, inefficient counterparts.
LLM training recipes and model initialization. The stacking method Gong et al. (2019); J. Reddi et al. (2023) employs a stage-wise training strategy that uses weights from initial layers to initialize later layers has been shown to be effective for LLM training both empirically Gong et al. (2019); J. Reddi et al. (2023); Du et al. (2024) and theoretically Agarwal et al. (2024). Knowledge distillation Hinton et al. (2015) has been very successful in training small LMs in some cases Turc et al. (2020); Sanh et al. (2019) the smaller student model is also initialized with teacher layers, though this is often done without clear explanation or intuition. Recent works in model initialization, such as Trockman & Kolter (2023), have studied synthetic attention patterns for initialization, primarily in vision settings. However, such methods have limited success in language models. Xu et al. (2024) use weight initialization for faster fine-tuning of vision models. In contrast, our proposed recipe focuses on creating smaller model by eliminating specific structural inefficiency in lazy layers. This distinction sets our work apart in terms of both objective and methodology.
6 Conclusion
In this paper, we identified a structural flaw in the attention mechanism of deep decoder-style LLMs, where many deeper layers tend to lose rank and converge into near rank-1 matrices. To address this, we propose $\mathsf{Inheritune}$, to train smaller models that inherits early blocks from a larger model and expands the architecture gradually, matching the performance of the reference model. We validated $\mathsf{Inheritune}$ on GPT-2 models of varying sizes, achieving efficient smaller models without performance loss on the OpenWebText and FineWeb_edu datasets.
Supplementary Materials
Contents
  • A: Extended Discussion on Attention Collapse
  • B: Baselines
  • C: Supplementary Experiments
  • D: Extended Discussion
  • E: Additional Experiments on Low Data Regime
  • F: Implementation Details
Appendix A Extended Discussion on Attention Collapse
A.1 Attention Mass Analysis
Dominant Single-Column Structure in Attention Matrices.
We further investigated the dominant structure of the rank-1 attention matrices and observed that, on an average, many of these matrices have their mass concentrated in a single column. This intrinsic structure can be viewed as a special case of rank-1 attention matrices. To quantify this, we computed the proportion of the matrix mass contributed by each column $j$ of $A(X)$ by computing $\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}$, where $A_{\cdot,j}$ denotes the $j$-th column of $A(X)$, $\|A_{\cdot,j}\|_{2}$ is the $\ell_{2}$-norm of that column, and $\|A(X)\|_{F}$ is the Frobenius norm of $A(X)$.
Next we determine the minimal number of columns required to capture $\eta$ proportion of the total mass, formally computed as;
$m^{*}=\min\left\{m\in\{1,2,\ldots,T\}\mid\sum_{j=1}^{m}\frac{\|A_{\cdot,j}\|_{2}^{2}}{\|A(X)\|_{F}^{2}}\geq\eta\right\},$
Here $\eta\in(0,1)$ represents the cumulative column mass threshold. In this work, we set $\eta=0.90$. A lower value of $m^{*}$ implies a stronger concentration of the attention matrix mass within fewer columns, reinforcing the phenomenon attention collapse. This analysis provides additional quantitative evidence highlighting the reduced representational capability of attention matrices in deeper transformer layers, further supporting the identification of lazy layers.
In Figure 8, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models. For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer. We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$. We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion. As shown in Figure 9, we performed a mass analysis on contemporary billion-parameter LLaMA models (Geng & Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers. This provides concrete evidence that the phenomenon persists across a broad range of architectures at the billion-parameter scale.
(a) GPT-2 Medium (355M) L=24, H=1024
(b) GPT-2 Large (770M) L=36, H=1280
(c) GPT-2 xLarge (1.5B) L=48, H=1600
Figure 8: In decoder-style LLMs, attention matrices in deeper layers often degenerate to near single column matrices, which is a special case of near rank-1. We compute $\text{AvgMass}^{(l)}$ (averaged over $N=100$ randomly selected sequences each with $T=100$ tokens) for each layer $l$ using the OpenWebText validation set. Our mass analysis of 24-layer GPT-2 medium, 36-layer GPT-2 large, and 48-layer GPT-2 xlarge models (L:layer, H:hidden size) reveals that attention matrices in many deeper layers collapse to single column matrices on an average.
(a) OpenLLaMA-3B
(b) OpenLLaMA-7B
(c) OpenLLaMA-13B
Figure 9: The overall mass of attention matrices in billion-scale LLMs, pre-trained on trillions of tokens, tends to concentrate in fewer columns. This phenomenon becomes increasingly pronounced as the model size grows. We computed attention matrices using 100 tokens from a random subset of RedPajama with 1B tokens. Next, we performed 100 runs and plotted the mean and standard deviation of the mass as a function of layers for our mass analysis, respectively. We followed the same procedure as discussed in Section 2. Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng & Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized. Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column. This observation aligns closely with our analysis in Figure 1.
A.2 Additional Analysis on Attention Rank Collapse.
To investigate the robustness of our analysis, we performed a rank analysis on GPT-2 models trained on OpenWebText (as shown in Figure 1) using a validation set from FineWeb-Edu—a new dataset on which the models were not trained (see Figure 10). It is clearly observed that attention consistently loses rank, often collapsing to single-column structures across all models, thereby reinforcing the robustness of our findings.
Next, we conducted a rank analysis on a contemporary LLaMA-3 model with 8B parameters. We compute $\text{Rank}^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}k^{*}_{n,h,l}$, where $N=100$ sequences are sampled from a subset of the C4 dataset (Raffel et al., 2019). As shown in Figure 11, we observe that nearly 50% of the attention heads (500 out of 1024 across all layers) are close to near rank-1, highlighted in red. This presents an interesting case: in very large modern architectures such as LLaMA-3 8B, while there may not be entire lazy layers, a substantial number of heads within many layers exhibit degeneracy.
A.3 Attention Pattern Visualization
To provide further evidence of lazy layers, we visualized attention patterns across various layers of a vanilla 24-layer GPT-2 medium model. Fig. 12 shows the attention patterns for both a vanilla 24-layer model trained from scratch and a 16-layer model trained using our proposed method, $\mathsf{Inheritune}$. Note just for the sake of better visualization we visualized full attention and not causal attention, in practice GPT-2 models computes causal attention. We computed these attention matrices using randomly selected strings from the val set of OpenWebText and took 40 tokens averaged over 3 runs.
In the 24-layer model trained from scratch (top row of Fig. 12), we observe a clear progression in attention patterns. The early layers (L4 and L7) exhibit structured patterns with a mix of local and global attention Gong et al. (2019); Beltagy et al. (2020); Chen et al. (2021). In contrast, the deeper layers (L20 and L22) display more uniform patterns, indicating a loss of focus. This uniformity is a hallmark of lazy layers, where the attention mechanism loses its ability to selectively focus on specific relevant tokens. In contrast, our 16-layer model trained with $\mathsf{Inheritune}$ (bottom row) demonstrates more focused and effective attention patterns, even in its later layers (L11 and L15). This striking difference suggests that our method makes model more attentive and addresses attention degeneration, potentially leading to more efficient models in compact size (also refer Figure 17 and Figure 18). Additionally, in Figure 13, we present attention patterns from a 48-layer GPT-2 xLarge model, comparing attention maps from a potent (non-lazy) and a lazy layer.
(a) Rank analysis of GPT-2 Medium
(b) Rank analysis of GPT-2 Large
(c) Rank analysis of GPT-2 XLarge
(d) Matrix mass analysis of GPT-2 Medium
(e) Matrix mass analysis of GPT-2 Large
(f) Matrix mass analysis of GPT-2 XLarge
Figure 10: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations. All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the Fineweb_edu validation set, following the same procedure described in Figure 1. This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
Figure 11: Rank analysis of LLaMA-3 8B reveals that nearly half of the attention heads exhibit rank collapse. We analyze the LLaMA-3 8B model, which contains 32 heads per layer (32 $\times$ 32), using the rank metric defined in Section 2. The results are visualized as a heatmap of head index vs. layer index. Potent (non-collapsed) heads are shown in varying shades of green, where higher intensity indicates higher rank, while rank-collapsed heads (near rank-1) are highlighted in red. Approximately 50% of all attention heads exhibit rank collapse, indicating widespread degeneracy.
L4 H2
L7 H3
L20 H2
L22 H4
24 layer GPT2 medium model trained from scratch.
L4 H2
L7 H3
L11 H2
L15 H1
Our 16-layer GPT2 medium model trained using $\mathsf{Inheritune}$.
Figure 12: $\mathsf{Inheritune}$ preserves effective attention patterns in smaller models. Comparison of attention patterns across layers (L) and heads (H) in two GPT2-medium models: (top) 24-layer model trained from scratch, (bottom) 16-layer model trained with $\mathsf{Inheritune}$. Attention maps are averaged over three randomly selected string, with 40 tokens each from the validation. Darker colors indicate higher attention scores. $\mathsf{Inheritune}$ maintains focused attention even in deeper layers, contrasting with the uniform patterns in the standard model’s deeper layers.
L8 H16
L8 H17
L8 H18
L8 H19
L8 H20
A Potent (non-lazy) layer of a pre-trained GPT2 xLarge 48 layer model.
L30 H16
L30 H17
L30 H18
L30 H19
L30 H20
A lazy layer of a pre-trained GPT2 xLarge 48 layer model.
Figure 13: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 xLarge model with 48 layers. The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
Appendix B Baselines
We compare $\mathsf{Inheritune}$ against several baselines. While some baseline methods are illustrated using GPT-2 Large or medium (for the knowledge distillation baseline) as an example, the same methodology is consistently applied across all model variants.
Baselines trained from scratch (rand init.) :
We compare our $\mathsf{Inheritune}$-derived model against much larger GPT-2 reference models trained from scratch for the same number of steps and similar-sized models trained from scratch for both the same and double the number of training steps.
Baselines trained with various model initialization and efficient training techniques.
Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various zeroshot model initialization and effcient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process for other models.
Stacking Gong et al. (2019); J. Reddi et al. (2023) is a model initialization and efficient (stage-wise) training recipe. We train a 9-layer GPT-2 large variant from scratch for 100K steps, then expanded the model to 18 layers by copying the weights from layers 0-8 to layers 9-17. Finally we re-trained this new 18-layer GPT-2 large variant, using stacking initialization for an additional 100K steps.
Hybrid stacking: Hybrid stacking is stacking but utilizes a large pre-trained reference model for initialization instead of using its own pre-trained weights. We took the weights of layers 0-8 from the reference 36-layer GPT-2 large model and expanded it to a 18-layer model by copying the weights to layers 0-17. We then trained this new 18-layer GPT-2 variant for 100K steps.
Half width: We initialized the baseline GPT-2 large variant across the width dimension and preserved the entire depth. We copied the weights of the first half the attention heads (0-9) and MLPs of the GPT-2 large reference model into baseline GPT-2 variant with half the width but all layers.
Baselines trained with Knowledge Distillation
As a baseline, we first apply logit-based knowledge distillation Hinton et al. (2015) to train a 16-layer GPT-2 medium variant (student) initialized randomly. For the second baseline, we use a DistillBERT-style approach Sanh et al. (2019), where the student model 0-11 layers are initialized with every alternate block of its teacher, and the remaining 4 blocks are initialized using layers 18, 19, 20, and 21 of the teacher. Both baselines are trained for 50K steps, using a vanilla 24-layer GPT-2 medium model as the teacher (our reference model).
Appendix C Supplementary Experiments
We provide additional training data and plots for our main results discussed in Section 4.1 as shown in Table 3 and Figure 14. In Figure 5 (also refer Table 3 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 14 (also refer Table 1) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
In Figure 15, we present the training curves of models trained during ablation as discussed in Section 4.2.
Knowledge Distillation
Recall we have already discussed distillation as a baseline in Section 4.1 and associated Figure 7. We perform an additional experiment in the same setting i.e. knowledge distillation as a baseline. Here we trained GPT-2 medium variants with 12 layers (half the number of a vanilla GPT-2 medium). We trained three models. First we distilled a 24-layer GPT-2 medium (teacher) to a 12-layer GPT-2 medium variant (student) and this student is initialized with all the alternate layers of the teacher. This setting is exactly same as discussed in DistillBERT Sanh et al. (2019). Next we trained two GPT-2 medium variants one from scratch (vanilla training) and the other with $\mathsf{Inheritune}$ recipe. Model trained with our recipe beats model trained with distillation. We defer a through investigation of distillation compared to $\mathsf{Inheritune}$ to future work.
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 12. Following up on our previous discussions in Figure 17 and Figure 18 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers compared to it’s larger counterpart trained form scratch. We performed rank analysis for Figure 17 utilizing vanilla 24-layer GPT-2 medium and our 16-layer GPT-2 variant trained using $\mathsf{Inheritune}$. Additionally, we performed rank analysis for Figure 18 with a vanilla 48-layer GPT2 xLarge and a 24-layer GPT2 xLarge variant trained using $\mathsf{Inheritune}$.
Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al. (2022). The vanishing gradients is caused when the norm of the gradients Bengio et al. (1994) are so small that it fails to generate meaningful back-propagation signal. Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
Models Layers Initialization Steps Pre-train Val loss ($\downarrow$)
GPT-2 Medium 24 rand init 100K 2.81
16 rand init 100K 2.86
12 Ours 100K 2.87
14 Ours 100K 2.84
Final Model $\longrightarrow$ 16 Ours 100K 2.81
GPT-2 Large 36 rand init 100K 2.85
18 rand init 100K 2.97
18 Ours 100K 2.80
GPT-2 xLarge 48 rand init 100K 2.65
24 rand init 100K 2.69
24 Ours 100K 2.64
Table 3: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts. All models are trained for 100K steps, and reported using validation loss on the OpenWebText validation set. Note: GPT-2 Large and xLarge use one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 14: Models derived using Inheritune outperform three zero-shot initialization and efficient training baselines in terms of final validation loss. Our models demonstrate better convergence and generalization compared to all baselines. We trained GPT-2 xlarge, GPT-2 large and GPT-2 medium variants on OpenWebText-9B for 100K steps using baseline model initialization and efficient training techniques and our $\mathsf{Inheritune}$ training recipe.
Figure 15: Full training curves of 16-layer GPT-2 variants trained during ablations. We analyze $\mathsf{Inheritune}$ approach while initializing some specific sub-modules in transformer blocks. Here, we initialize each transformer block of a 16-layer GPT-2 medium variant with three different configurations. First, we separately initialize attention and MLPs (FFNs) submodules; second, we initialize the attention and MLP weights while randomly initializing the layer norms. Finally, we perform $\mathsf{Inheritune}$-initialize only the attention and MLP weights with all the respective layer norms.
Figure 16: A 12-Layer GPT-2 medium variant derived using Inheritune converges faster and generalizes better than a same-sized models trained from scratch and with Logit-based distillation with teacher initialization baseline. Three 12-layer GPT-2 medium variants were trained: (1) a distilled model initialized with alternate layers from a 24-layer GPT-2 medium teacher, following the DistillBERT setup Sanh et al. (2019); (2) a model trained from scratch (vanilla training); and (3) a model trained using the $\mathsf{Inheritune}$ recipe. The model trained with $\mathsf{Inheritune}$ outperforms both the distillation-based model and the one trained from scratch, demonstrating the effectiveness of our approach.
(a) Vanilla GPT2-Medium 24 layers
(b) Our GPT2-Medium 16 layer variant trained with $\mathsf{Inheritune}$
Figure 17: Rank collapse in deeper layers and its mitigation through $\mathsf{Inheritune}$. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 24-layer GPT2 medium model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 3 out of the last 12 later layers exhibit rank-1 attention matrices (mean rank accross all the 100 runs). (b) Our 16-layer GPT2 medium variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, none of the later layers in our 16-layer variant exhibit rank-1 attention matrices.
(a) Vanilla GPT2-xLarge 48 layers
(b) Our GPT2-xLarge 24 layer variant trained with $\mathsf{Inheritune}$
Figure 18: Rank collapse worsens for larger LLMs, $\mathsf{Inheritune}$ helps to mitigate rank collapse. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Fig. 1 (a) Analysis of a 48-layer GPT2 xLarge model reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs). (b) Our 24-layer GPT2 xLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across all layers, highlighting the effectiveness of our approach. Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
Appendix D Extended Discussion
D.1 Discussion about Attention Sink
The term "attention sink" Xiao et al. (2024) refers to the phenomenon where the first token in a sequence receives disproportionately high attention scores compared to other tokens in the attention maps. While there is some connection with $\mathsf{Inheritune}$, as we have also observed that many attention matrices are not only rank-1 but also single-column (with all attention scores concentrated on the first token), this connection has not been explicitly established in Xiao et al. (2024) with respect to rank-1 behavior or poor training of later layers.
In contrast, as illustrated in Figure 1, we compute the maximum rank of all attention matrices within a layer. For instance, consider a layer where only 2 out of 5 attention heads exhibit attention sink behavior. This does not make the layer lazy, as attention is computed as a concatenation of activations across all heads. A lazy layer, however, has all 5 out of 5 attention heads fully degenerated, with their attention matrices being rank-1. We have provided evidence that such lazy layers are indicative of poorly trained layers.
Layers Initialization Avg max ranks Val Loss ($\downarrow$)
4 rand n/a 3.25
4 1-4 layers from vanilla GPT2 8.40 3.22
4 5-8 layers from vanilla GPT2 9.48 3.19
4 9-12 layers (lazy layers) from GPT2 1.22 3.23
Table 4: Impact of initialization strategies on GPT2-small variants. We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps. Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText. Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer. The training plots and rank analysis are provided in Figure 2.
Models Recipe Layers ARCE PIQA SciQ Hellaswag Lambada Average
(acc) (acc) (acc) (acc norm) (acc)
GPT-2 Medium rand init 24 51.05 61.81 74.8 30.79 20.28 47.74
rand init 16 49.92 61.92 73.3 29.56 19.54 46.84
Ours 16 51.26 61.81 73.8 30.55 23 48.08
GPT-2 Large rand init 32 52.48 64.58 75.3 32.65 22.2 49.44
rand init 16 50.34 63.11 75 30.86 21.56 48.17
Ours 16 52.9 63.55 76.1 32.14 24.06 49.75
Table 5: Models trained with Inheritune outperforms both their larger and same-size counterparts trained from scratch on average zero-shot downstream performance. For evaluation we choose accuracy (acc) and normalized accuracy (acc norm) metrics following Open LLM leaderboard Beeching et al. (2023). All the models are trained with FineWeb_edu.
Appendix E Developing a 1.5B Small Base LM in a Low Data Regime with Inheritune
Figure 19: Performance of our 1.5B base LM derived using 1B data with $\mathsf{Inheritune}$ on an average of 9 different datasets (left) and MMLU benchmark (right) that evaluates commonsense, truthfulness, natural language inference and language understanding. We compare our model’s performance with reference model-OpenLLamA-3B (2x size), other small base LMs of size 1B-2B parameters such as MPT-1.3B, OPT-1.3B, Pythia-1.4B (pre-trained from scratch) and ShearLLaMA-1.5B (pruned and continually trained using existing large base LM).
In this section, we aim to investigate the efficacy of $\mathsf{Inheritune}$ in a data and compute-constrained setting. We train a 1.5B parameter small base LM with only 1B tokens using a 3B parameter base LM on a single GPU (A6000) for less than half a day.
We assume the existence of a pre-trained reference model $\mathcal{M}_{\text{ref}}$, comprising $k$ layers, represented by $\mathsf{W}_{\text{ref}}=\{\mathsf{W}_{0},\mathsf{W}_{1},\ldots,\mathsf{W}_{k-1}\}$ trained with $\mathcal{D}_{\text{train}}$. However, this full training data is unavailable, and we only have a random tiny subset $\hat{\mathcal{D}}_{\text{train}}\sim\mathcal{D}_{\text{train}}$. We use OpenLLaMA-3B version 1 as the reference model pre-trained with 1T tokens from the RedPajama V1 dataset, which contains data from various domains such as common crawl, C4, Wikipedia, books, arXiv papers, GitHub, and Stack Exchange. We take 1B randomly sampled tokens [^3] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T-Sample from the RedPajama dataset.
Training recipe.
To adapt $\mathsf{Inheritune}$ for this new setting, we perform step 1 and step 2 in Algorithm 1 without growing the model (i.e., we skip step 3). We use the first $n=13$ layers from our $k=26$ layer reference model. We call our small base LM Ours-1.5B(#tokens). We train our model with data repetition for eight epochs (each epoch uses all the 1B tokens) with a batch size of 131K tokens per batch. We use 1 A6000 GPU for less than half a day of training. The choice of training epochs is based on the analysis provided later in this paper (refer to Figure 22). We use the lit-gpt framework for training all small base LMs discussed in this paper. Further discussions on the training hyper-parameters can be found in the next Section.
Baseline models and evaluation.
We choose similarly sized (1-2B parameter) small base LMs trained with the RedPajama dataset and the reference base LM as primary baselines, as the quality of the pre-training data plays a key role in model development. We also include models OPT-1.3B Zhang et al. (2022) and Pythia-1.3B Biderman et al. (2023) as these models are pre-trained with a dataset similar to the RedPajama dataset. Table 7 lists the baseline models with their pre-training data.
In this study, we use few-shot accuracy, particularly 0-shot and 5-shot accuracy, on ten different downstream tasks to measure the quality of our 1.5B base LM. This evaluation of pre-trained base LLMs has been done in several prior works. Our evaluation methodology categorizes downstream tasks across four distinct genres: commonsense reasoning, natural language understanding, factuality, and natural language inference. We perform 0-shot evaluation for PIQA Bisk et al. (2020), BOOLQ Clark et al. (2019), WINOGRANDE Sakaguchi et al. (2020), WINOGRAD Kocijan et al. (2020), LOGIQA Liu et al. (2020), TruthfulQA Lin et al. (2022), MNLI Bowman et al. (2015), QNLI Wang et al. (2018) and WNLI Wang et al. (2018) datasets. Next, we perform a 5-shot evaluation on the massive multitask language understanding benchmark (MMLU) Hendrycks et al. (2020). We use the lm eval harness framework Gao et al. (2024) for the entire evaluation.
Model Commonsense Reasoning
Name (# train tokens) Reference Winograd PIQA Boolq WinoGrande Logiqa
OpenLLaMA-3B (1T) n/a 63.46 74.97 67.18 62.27 28.4
OPT-1.3B (300B) n/a 38.46 71.82 57.83 59.51 27.04
Pythia-1.4B (300B) n/a 36.54 70.89 63.12 56.99 27.65
MPT-1.3B (200B) n/a 63.46 71.44 50.89 58.09 28.26
Sheared LLaMA-1.3B (50B) LLaMA2-7B 36.54 73.45 62.02 58.17 27.34
Ours-1.5B (1B) OpenLLaMA-3B 50.96 56.47 61.68 51.69 25.19
Model Lang. Understanding & Inference Factuality
Name ( # train tokens) Reference MMLU(5) WNLI QNLI MNLI TruthfulQA
OpenLLaMA-3B (1T) n/a 27.21 50.7 51.3 37.3 35
OPT-1.3B (300B) n/a 24.96 42.25 51.29 35.82 38.67
Pythia-1.4B (300B) n/a 25.56 53.52 49.48 32.76 38.66
MPT-1.3B (200B) n/a 25.82 40.85 50.52 35.93 38.68
Sheared LLaMA-1.3B (50B) LLaMA2-7B 25.71 49.3 50.98 37.94 37.14
Ours-1.5B (1B) OpenLLaMA-3B 25.67 43.66 49.41 34.42 48.61
Table 6: Our 1.5B model achieves performance comparable to baseline models despite being trained with fewer tokens. Comparison of our target model ($\mathcal{M}_{\text{tgt}}$) derived using $\mathsf{Inheritune}$ with the reference model ($\mathcal{M}_{\text{ref}}$) and other baseline models of similar size when pre-trained from scratch and pre-trained with inherited weights and pruning. Although trained with fewer tokens, our model achieves performance comparable to the baseline models. We have highlighted all the scores in bold where our 1.5B model achieves at least 90% of the score compared to the reference LM or outperforms at least two of the baseline similar-size LMs. All the tasks are evaluated using 0-shot except MMLU, which is 5-shot. The models against which n/a is mentioned are trained from scratch.
E.1 Main Results in Low Data Regime
Table 6 presents a detailed performance evaluation across various tasks. Our 1.5B model, developed using $\mathsf{Inheritune}$, excels in 7 out of 10 individual tasks. It achieves a score of 90% or higher compared to the reference language model, which is twice its size and trained with 1000 times more data, or it outperforms at least two other base LMs of similar size trained with 50-300 times more data. Favorable scores are highlighted in bold.
Next, we compare our small LM with the MPT-1.3B [^4] https://huggingface.co/mosaicml/mpt-1b-redpajama-200b model trained from scratch with 200B tokens of RedPajama dataset and find that we match 97% accuracy in all nine downstream tasks and the MMLU (5-shot) score. Additionally, we compare with OPT-1.3B and Pythia-1.3B models, showing that we outperform both in the MMLU (5-shot) score and perform comparably on the other nine datasets. This comparison illustrates that having a large reference base LM and a subset of its pre-training data allows the inherited target size base LM to be trained remarkably more sample-efficiently than training from scratch. Extended discussions on comparisons with the ShearedLLaMa model, generated by pruning and continual training from LLaMA2-7B, are provided in the supplementary materials.
Ablation of Inheritune Across Different Model Sizes with 1B Tokens.
In the previous section, we considered a single choice of $n=k/2$, i.e., half the layers, for the size of the smaller model. Here, we investigate $\mathsf{Inheritune}$ with different choices of $n$, but the same 1B token dataset). All models use OpenLLAMA-3B as the large pre-trained reference model, with consistent training hyperparameters, changing only the choice of $n$.
We developed eight different submodels with $n=\{4,6,8,10,13,16,18,20\}$. Figure 20 shows the MMLU (5-shot) score as a function of $n$. As expected, the trend line is positive-sloping. The submodel with 20 layers slightly decreases performance, potentially due to data overfitting as the model size increases. The training details for all these submodels are consistent with the target 1.5B small base LM and are detailed in the appendix. A more comprehensive investigation on the choice of $n$—including varying both $n$ and the number of training tokens jointly and evaluating a broader set of tasks is left for future work.
Figure 20: Inheritune scales across multiple different model sizes. Utilizing the OpenLLaMA-3B as a reference large base LM, demonstrates that multiple performant small base LMs of target size can be crafted using $\mathsf{Inheritune}$ with just 1B training tokens. The MMLU (5-shot) as a function of the number of submodels.
E.2 Additional analysis with larger reference LMs and 50B data
We further analyze $\mathsf{Inheritune}$ to see the impact of it’s performance when more tokens are available. Initially for the main results we limited ourselves to 1B (i.e. 0.1%) tokens from the 1T pre-training data, here we use a 50B subset (i.e. 5%) of the pre-train data. Moreover we also extend this study to include larger base LMs of 7B parameters as reference models, employing OpenLLaMA-7B and LLaMA2-7B as reference models. For the purpose of this study we do not repeat the tokens from our 50B subset. As shown in Figure 21, we observe that there is clear improvement in overall MMLU (5-shot) score with more data. Additionally it is interesting to see that 1.5B (or 1.6B models) developed with $\mathsf{Inheritune}$ using larger reference models show even greater improvements when fed with 50B subset of non repetitive data (i.e fresh tokens). We present a Table 9 using Figure 21 to show the best MMLU (5-shot) scores achieved using different reference LMs. For developing our small base LMs using larger reference LMs we use $n$=7 (i.e. 7 layers). The training details are discussed in the following section.
Model Training Data (# tokens)
OpenLLaMA-3B v1(ref) RedPajama(1T)
Ours-1.5B* RedPajama (1B)
Shear-LLaMA-1.3B* RedPajama(50B)
MPT-1.3B RedPajama(200B)
Pythia-1.4B The Pile(300B)
OPT-1.3B Custom data(300B)
Table 7: Comparison of training data across baseline models. Overview of reference and baseline models, including their pre-training datasets and the number of tokens used during training. Note the significant variation in training data size, ranging from 1B to 1T tokens.
Models (# train tokens) GPU Count GPU Type Time (# days)
MPT-1.3B (200B) 440 A100 half
Pythia-1.4B (300B) 64 A100 4.6
TinyLLaMA-1.1B (3T) 16 A100 90
OPT-1.3B (300B) 992 A100
Sheared LLaMA-1.3B (50B) 16 A100
OpenLLaMA-3B (1T) 256 TPU v4 10
Our-1.5B (1B) 1 A6000 ~half
Table 8: Computational efficiency of Inheritune versus baseline models. Comparison of pre-training compute requirements for publicly available small base LMs and our $\mathsf{Inheritune}$-derived model. Metrics include GPU count, GPU type, and training duration, highlighting Inheritune’s significant reduction in computational resources.
Figure 21: Impact of reference model choice on Inheritune performance. MMLU (5-shot) scores for 1.5B base LMs derived using $\mathsf{Inheritune}$, trained on 50B unique tokens. Comparison across three reference models: OpenLLaMA-7B, LLaMA2-7B, and OpenLLaMA-3B. Results demonstrate Inheritune’s effectiveness with various large language models as references.
Figure 22: Performance of our 1.5B base LM derived using $\mathsf{Inheritune}$ based on existing OpenLLaMA-3B base model. Here we use 1B tokens and perform data repetition (epochs) during training. We further evaluate our model on an average of 9 different datasets (left) and MMLU benchmark (right).
Model (# tokens), ref MMLU(5)
Ours-1.6B (1B), LLaMA2-7B 24.27
Ours-1.5B (1B), OpenLLaMA-3B 25.67
Ours-1.5B (50B), OpenLLaMA-3B 25.71
Ours-1.6B (50B), LLaMA2-7B 26.07
Ours-1.6B (50B), OpenLLaMA-7B 26.72
Table 9: Performance comparison of models on the MMLU (5-shot) task. Our models, even when trained with fewer tokens, show competitive performance compared to benchmarks. We have highlighted the best MMLU 5-shot score in bold.
Model (# tokens) Data type MMLU (5-shot)
Ours-1.5B (1B) 10 epochs 24.95
Ours-1.5B (50B) 10B fresh 23.62
Ours-1.5B (1B) 20 epochs 25.46
Ours-1.5B (50B) 20B fresh 24.96
Table 10: MMLU (5-shot) scores of Our-1.5B small base LM derived using 1B data for multiple data repetition–10 epochs and 20 epochs compared to the same model trained without data repetition for 10B and 20B fresh tokens. We derive all the variants of Our-1.5B small base using $\mathsf{Inheritune}$ with OpenLLaMA-3B as reference model. The models featured in this table correspond to those discussed in Figures 21 and 22.
Ablations with number of epochs.
We ran ablations (refer Figure 22) to choose the total number of epochs (multiple passes over the data) and observe that repetition when training our 1.5B (or 1.6B) LM is helpful particularly for MMLU. We also observe that the for an average of all the 9 other datasets (i.e. except MMLU) peaks it’s performance at 5 epochs and then deteriorates. Some prior works have studied this phenomenon that the scaling of downstream tasks with data is not always linear Biderman et al. (2023).
To repeat or not to repeat the tokens.
Next we tackle the question – whether one should re-use 1B tokens for multiple epochs or use the same number of fresh tokens? Some prior works have recommended that if you have a reasonably large size dataset one can repeat it upto 4 epochs Muennighoff et al. (2023). In our study we observe that one can safely re-use 1B tokens upto 10-20 epochs as shown in Table 10. We emphasis that this phenomenon needs a through investigation in itself and we defer this to future work. The models discussed in Table are saved checkpoints during a single training run and not the final model unless otherwise specified.
E.3 Implications of Low Data Regime
In this section, we discuss some of the key implications of our work in low data regime.
Cheap and easy development of small base LMs.
Pre-training a small base LM of 1-2B parameters from scratch is extremely expensive. For instance mpt-1.3B base LM is pre-trained with 440 A100 GPUs for half a day, while the Pythia-1.4B base LM Biderman et al. (2023) utilized 64 A100-40GB GPUs for 4.6 days. Similarly, TinyLLaMA-1.1B model Peiyuan Zhang & Lu (2023) was pre-trained using 16 A100 GPUs for 3 months. Our 1.5B (1B data variant) LM shows competitive performance despite being trained with 1 A6000 GPU for less than 12 hours. The computational details are provided in Table 8, comparing the training resources of the baseline models listed in this paper. Typically small base LMs are finetuned for a specific task before deployment and are not used in it’s base form. With $\mathsf{Inheritune}$ we present a really easy and cheap way for developing a small base LM to be later finetuned before deployment.
Naive baseline for pre-training a scaled down variant of large base LMs.
Typically small variants of large base LMs are pre-trained using the same pre-training data Peiyuan Zhang & Lu (2023); Groeneveld et al. (2024). Our recipe introduces a new perspective of identifying sufficient depth without losing any generalization on the held out validation set. Next, we also show that even with a small fraction of pre-train data (randomly sampled) and few initial layers of the large base LM one can develop a small base LM. Therefore our $\mathsf{Inheritune}$ recipe has the potential to become the naive baseline for any pre-training pipeline aiming to develop a smaller variant of a large base LM.
Appendix F Implementation Details
F.1 Training details of GPT-2 models
For our main experiments, we focused on three sizes of GPT-2 models Radford et al. (2019): the vanilla GPT-2 xlarge with 1.5B parameters, GPT-2 large with 770M parameters and the vanilla GPT-2 medium with 355M parameters. We developed several variants of these models by adjusting the number of layers and hidden size. We trained all GPT-2 models with data repetition while using OpenWebText dataset, the trainset has 9B tokens and the validation set has 4.4M tokens. The key architectural configurations for the reference models, our models, and baseline models discussed in this paper are summarized in Table 11.
For all training runs, we used GELU activations, disabled bias terms, and removed dropout, following the nanoGPT codebase and Liu et al. (2023). We employed the AdamW optimizer with $\beta_{1}=0.90$ and $\beta_{2}=0.95$. The GPT-2 models were trained on a single node with 3 A100 GPUs (each with 40 GB of memory) using distributed data parallelism and gradient accumulation. In line with Liu et al. (2023), we scaled the attention logits inversely to the layer index across all GPT-2 models. Most hyperparameters were adapted from Liu et al. (2023), with key details provided below.
Hyper-parameter details of GPT-2 Medium and variants.
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 large and variants.
  • Batch size: 16K tokens
  • Learning rate: $2\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of GPT-2 xlarge and variants.
  • Batch size: 16K tokens
  • Learning rate: $1.5\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decayed to $1\times 10^{-5}$,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 100K
Hyper-parameter details of knowledge distillation training.
We use the below loss for as our distillation based training loss. The validation loss is the student_loss.
[…]
  • Model: 16-layer and 12-layer GPT-2 medium variants
  • Softmax temperature: 1
  • $\alpha$: 0.6
  • Batch size: 50K tokens
  • Learning rate: $3\times 10^{-4}$,
  • Warmup steps: 2K,
  • Scheduler type: cosine decay to $\frac{1}{10}$ of max learning rate,
  • Weight decay: 0.1,
  • Gradient clipping value: 1,
  • Total training steps: 50K
Models Layers Hidden Size Heads Variant
[…]
GPT2-large(770M) 36 1280 20 Original
GPT2-large(680M) 32 1280 20 Original
GPT2-medium(355M) 24 1024 16 Original
[…]
GPT2-medium 16 512 8 half width
[…]
GPT2-large 18 1280 20 Ours
GPT2-large 16 1280 20 Ours
GPT2-medium 16 1024 16 Ours
Table 11: Overview of all the GPT2 models used in this study and their architectural configurations. The model configurations of stacking and hybrid stacking are same as our variants.
F.2 Training details of 1.5B OpenLLaMA model
Small base LMs trained with 1B data
We present our main results with Our-1.5B model trained with an existing OpenLLaMA version 1 Geng & Liu (2023) and 1 B tokens randomly sampled from 1T redpajama version1 data. The hyper-parameters related to this model is provided below. It is important to note that our claim that we only use 1 GPU for less than 12 hours to train Our-1.5 B model is specific to models derived using $\mathsf{Inheritune}$ with 1B data. Next we also train multiple sub-models as shown in Figure 20 the training details remains consistent with that of the initial model discussed earlier. However we observe that increasing the number of layers in a sub-model also increase the training time.
Hyper-parameter details of our 1.5B base LM derived using OpenLLaMA-3B as refernce LM:
  • Training tokens: 1B
  • Training epochs: 8
  • Training steps: 64K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm up steps: 1000
  • Batch size: 131K
  • GPU count: 1
  • GPU type: A6000
  • GPU hours: $\sim$ 8 hours
  • GPU hours/epoch: $\sim$ 54 minutes
Training details of small base LMs with 50B data.
We also trained our 1.5B model with larger subsets of data as shown in Figure 21.It is important to note that all the intermediate tokens until 50B are intermediate checkpoints of a single training run. Some of the key hyper-parameters of our training runs are discussed below. We have also trained three variants of small base LMs utilizing 3 different reference base LMs namely OpenLLaMA-3B, OpenLLaMA-7B and LLaMA2-7B. For target LMs developed with OpenLLaMA-3B we use $n$=13 i.e. 13 layers. For target LMs developed using reference LMs of 7B parameters we use $n$=7 i.e. 7 layers. The training hyper-parameters remains consistent across all the models trained with 50B subset of the pre-train data.
Training hyper-parameters of our target 1.5B and 1.6B small base LMs:
  • Training tokens: 50B
  • Training epochs: $\sim$1
  • Training steps: 191K
  • Learning rate: $3\times 10^{-4}$
  • Scheduler: Cosine
  • Weight decay: 0.1
  • Optimizer: AdamW
  • Warm-up steps: 1000
  • Batch size: 131K tokens
  • GPU count: 1
  • GPU type: A100
  • GPU hours: $\sim$18 hours